### v1

# Cuspidal edges on focal surfaces of regular surfaces with parabolic points

Abstract.
We investigate geometric invariants of cuspidal edges on focal surfaces of regular surface. In particular, we shall clarify the sign of the singular curvature at a cuspidal edge on a focal surface using singularities of parallel surface of a given surface satisfying certain conditions.
Key words and phrases:
parallel surface, focal surface, singularity, cuspidal lips, cuspidal beaks, Gaussian curvature
2020 Mathematics Subject Classification:
53A05, 53A55, 57R45
The author was partially supported by JSPS KAKENHI Grant Numbers JP22K13914 and JP22KK0034.
1. Introduction
For a regular surface in Euclidean $3$-space $\mathbb{R}^{3}$, its parallels and focal surfaces can be examined. While the initial surface is non-singular, its parallels and focal surfaces generally have singularities ([10]). From the viewpoint of singularity theory of smooth mappings, singularities that appear on parallel surfaces are Legendrian singularities, while those on focal surfaces are Lagrangian singularities ([1, 2, 10]). In both cases, a cuspidal edge is known to be a generic singularity ([1]). It has been established that there is a relationship between singularities of parallels and focal surfaces of a regular surface ([8, 26]). See Table 1 for example. We remark that parallel surfaces are fronts ([1, 2, 10, 28]).
singularity of parallel surface singularity of focal surface
cuspidal edge regular
swallowtail cuspidal edge
cuspidal lips cuspidal edge
cuspidal beaks cuspidal edge
Table 1. Relation between singularities (see Proposition 4.1).
Conversely, there are numerous studies on surface with cuspidal edge singularities from the perspective of differential geometry (cf. [7, 16, 14, 19, 22, 24, 25]). At a cuspidal edge, several geometric invariants are introduced ([16, 14, 22]). In particular, geometric invariants called the limiting normal curvature and the singular curvature play a crucial role. It has been established that the limiting normal curvature is associated with the boundedness of the Gaussian curvature at a cuspidal edge ([22, 16]). While the singular curvature relates to convexity or concavity of cuspidal edge. More precisely, the following fact is known ([22, Theorem 3.1]): The Gaussian curvature is bounded near a cuspidal edge if and only if the limiting normal curvature vanishes along the edge. Moreover, if the Gaussian curvature takes non-negative values near a cuspidal edge, then the singular curvature is non-positive at the cuspidal edge. For a pseudo-spherical front, that is, a surface with certain singularities whose Gaussian curvature is negative constant on the set of regular points, the Gaussian curvature of its focal surface is bounded (cf. (A.7)). In addition, we see the following assertions ([27]): If the initial pseudo-spherical front has a cuspidal lips (resp. cuspidal beaks), then the singular curvature of the focal surface is positive (resp. negative) at the corresponding cuspidal edge. Thus it is natural to ask that how singularities of fronts (parallel surfaces) relate to geometric properties of cuspidal edges on focal surfaces.
In this paper, we investigate the limiting normal curvature and the singular curvature at cuspidal edges of focal surfaces of regular surfaces. In particular, we deal with a surface whose Gauss map has singularity at some point. In that case, the Gaussian curvature of parallel surfaces satisfies certain bounded condition (see Propositions 3.3 and 3.4). Moreover, the limiting normal curvature defined at a corank one singular point on the parallel surface vanishes in the case (Proposition 3.5). Comparing conditions for singularities of parallel surfaces and focal surfaces, it holds that the focal surface is a cuspidal edge if the parallel surface is a swallowtail, a cuspidal edge or a cuspidal beaks (Proposition 4.1). For a curvature line parametrized surfaces, we give explicit representations for the limiting normal curvature and the singular curvature at a cuspidal edge of the focal surface (Propositions 4.3 and 4.5). In particular for the singular curvature, we shall show that relation between the sign of the singular curvature at a cuspidal edge of the focal surface and singularities of parallel surface (Theorem 4.6).
Acknowledgments
The author thanks Kentaro Saji and Runa Shimada for fruitful discussions and valuable comments.
2. Preliminaries
2.1. Surfaces in Euclidean $3$-space
We here review some fundamental concepts from differential geometry of immersed surfaces in Euclidean $3$-space $\mathbb{R}^{3}$. For details, see [6, 9, 21].
2.1.1. Surfaces parametrized by curvature line coordinates
Let $U\subset\mathbb{R}^{2}$ be a domain. Let $f\colon U\to\mathbb{R}^{3}$ be an immersion and $\nu$ its unit normal vector. Then we call the image of $f$ or $f$ itself a surface. We assume that $f$ is a surface without umbilic point in the following. Then we can take a curvature line coordinate system $(u,v)$ on $U$. In this case, it holds that
$\left\langle{f_{u}},{f_{v}}\right\rangle=\left\langle{\nu_{u}},{f_{v}}\right\rangle=\left\langle{\nu_{v}},{f_{u}}\right\rangle=0,$
where $\left\langle{\cdot},{\cdot}\right\rangle$ is the canonical inner product on $\mathbb{R}^{3}$, $(\cdot)_{u}=\partial(\cdot)/\partial u$ and $(\cdot)_{v}=\partial(\cdot)/\partial v$. Moreover, the first and second fundamental forms of $f$ can be written as
$\operatorname{I}=Edu^{2}+Gdv^{2},\quad\operatorname{II}=Ldu^{2}+Ndv^{2},$
where
$E=\left\langle{f_{u}},{f_{u}}\right\rangle,\quad G=\left\langle{f_{v}},{f_{v}}\right\rangle,\quad L=-\left\langle{f_{u}},{\nu_{u}}\right\rangle,\quad N=-\left\langle{f_{v}},{\nu_{v}}\right\rangle.$
Then the Gauss and Weingarten formulas are as follows:
(2.1) $f_{uu}$ $=\Gamma^{1}_{11}f_{u}+\Gamma^{2}_{11}f_{v}+L\nu=\frac{E_{u}}{2E}f_{u}-\frac{E_{v}}{2G}f_{v}+L\nu,$
$f_{uv}$ $=f_{vu}=\Gamma^{1}_{12}f_{u}+\Gamma^{2}_{12}f_{v}=\frac{E_{v}}{2E}f_{u}+\frac{G_{u}}{2G}f_{v},$
$f_{vv}$ $=\Gamma^{1}_{22}f_{u}+\Gamma^{2}_{22}f_{v}+N\nu=-\frac{G_{u}}{2E}f_{u}+\frac{G_{v}}{2G}f_{v}+N\nu,$
$\nu_{u}$ $=-\kappa_{1}f_{u},\quad\nu_{v}=-\kappa_{2}f_{v},$
where $\kappa_{1}$ and $\kappa_{2}$ are principal curvatures of $f$ given by
$\kappa_{1}=\frac{L}{E},\quad\kappa_{2}=\frac{N}{G}.$
Moreover, Codazzi equations can be written as
(2.2) $L_{v}=\frac{1}{2}(\kappa_{1}+\kappa_{2})E_{v},\quad N_{u}=\frac{1}{2}(\kappa_{1}+\kappa_{2})G_{u}.$
Lemma 2.1.
Under the above setting, we have
$\frac{E_{v}}{2E}=\Gamma^{1}_{12}=\frac{(\kappa_{1})_{v}}{\kappa_{2}-\kappa_{1}},\quad\frac{G_{u}}{2G}=\Gamma^{2}_{12}=\frac{(\kappa_{2})_{u}}{\kappa_{1}-\kappa_{2}}.$
Proof.
We consider $\kappa_{1}$. By the definition, we have
$(\kappa_{1})_{v}=\frac{L_{v}}{E}-\frac{LE_{v}}{E^{2}}=\frac{L_{v}}{E}-\kappa_{1}\frac{E_{v}}{E}.$
By Codazzi equations (2.2), it holds that
$(\kappa_{1})_{v}=\frac{E_{v}}{2E}(\kappa_{1}+\kappa_{2})-\kappa_{1}\frac{E_{v}}{E}=\frac{E_{v}}{2E}(\kappa_{2}-\kappa_{1}).$
Thus we obtain the assertion for $\kappa_{1}$. For $\kappa_{2}$, one can show in a similar way. ∎
By (2.1) and Lemma 2.1, it holds that
$f_{uv}=-\frac{(\kappa_{1})_{v}}{\kappa_{1}-\kappa_{2}}f_{u}+\frac{(\kappa_{2})_{u}}{\kappa_{1}-\kappa_{2}}f_{v}.$
Definition 2.2.
For a surface $f\colon U\to\mathbb{R}^{3}$ parametrized by curvature line coordinate system $(u,v)$,
  • a point $p\in U$ is said to be a $\partial_{u}$-ridge point (resp. $\partial_{v}$-ridge point) if $(\kappa_{1})_{u}(p)=0$ (resp. $(\kappa_{2})_{v}(p)=0$);
  • a point $p\in U$ is said to be a $\partial_{u}$-sub-parabolic point (resp. $\partial_{v}$-sub-parabolic point) if $(\kappa_{2})_{u}(p)=0$ (resp. $(\kappa_{1})_{v}(p)=0$).
Ridge points on a surface are originally introduced by Porteous to investigate singularities of distance squared functions on the surface ([20]). For details of ridge and sub-parabolic points on a surface, see [4, 3, 20, 21, 18].
2.1.2. Curvature line frame
Let $f\colon U\to\mathbb{R}^{3}$ be a regular surface without umbilic point. Let $\nu$ be a unit normal vector to $f$. We then take a curvature line coordinate system $(u,v)$ on $U$. We define maps $\bm{e}_{i}\colon U\to\mathbb{R}^{3}$ ($i=1,2$) by
(2.3) $\bm{e}_{1}=\frac{f_{u}}{|f_{u}|},\quad\bm{e}_{2}=\frac{f_{v}}{|f_{v}|}.$
Then it holds that
$\left\langle{\bm{e}_{i}},{\bm{e}_{j}}\right\rangle=\delta_{ij},\quad\nu=\bm{e}_{1}\times\bm{e}_{2},$
where $\delta_{ij}$ is the Kronecker delta. Thus the triplet $\{\bm{e}_{1},\bm{e}_{2},\nu\}$ forms an orthonormal frame along $f$. We call this frame the curvature line frame along $f$.
Lemma 2.3 (cf. [5]).
With the above condition, we have
$\begin{pmatrix}\bm{e}_{1}\\ \bm{e}_{2}\\ \nu\end{pmatrix}_{u}=\begin{pmatrix}0&x_{1}&x_{2}\\ -x_{1}&0&0\\ -x_{2}&0&0\end{pmatrix}\begin{pmatrix}\bm{e}_{1}\\ \bm{e}_{2}\\ \nu\end{pmatrix},\quad\begin{pmatrix}\bm{e}_{1}\\ \bm{e}_{2}\\ \nu\end{pmatrix}_{v}=\begin{pmatrix}0&y_{1}&0\\ -y_{1}&0&y_{3}\\ 0&-y_{3}&0\end{pmatrix}\begin{pmatrix}\bm{e}_{1}\\ \bm{e}_{2}\\ \nu\end{pmatrix}$,
where
$x_{1}=\frac{(\kappa_{1})_{v}|f_{u}|}{(\kappa_{1}-\kappa_{2})|f_{v}|},\quad x_{2}=\kappa_{1}|f_{u}|,\quad y_{1}=\frac{(\kappa_{2})_{u}|f_{v}|}{(\kappa_{1}-\kappa_{2})|f_{u}|},\quad y_{3}=\kappa_{2}|f_{v}|.$
Proof.
For $\nu$, it holds that
$\nu_{u}=-\kappa_{1}f_{u}=-\kappa_{1}|f_{u}|\bm{e}_{1},\quad\nu_{v}=-\kappa_{2}f_{v}=-\kappa_{2}|f_{v}|\bm{e}_{2}$
by Weingarten formulas (2.1). By (2.3) and Lemma 2.1, $(\bm{e}_{1})_{u}$ and $(\bm{e}_{2})_{u}$ can be calculated as
$(\bm{e}_{1})_{u}$ $=\frac{f_{uu}}{|f_{u}|}-\frac{f_{u}(|f_{u}|)_{u}}{|f_{u}|^{2}}=\frac{1}{|f_{u}|}\left(\Gamma^{1}_{11}f_{u}+\Gamma^{2}_{11}f_{v}+L\nu\right)-\Gamma^{1}_{11}\bm{e}_{1}$
$=-\frac{E_{v}}{2G|f_{u}|}f_{v}+\frac{L}{|f_{u}|}\nu=\frac{(\kappa_{1})_{v}|f_{u}|}{(\kappa_{1}-\kappa_{2})|f_{v}|}\bm{e}_{2}+\kappa_{1}|f_{u}|\nu,$
$(\bm{e}_{2})_{u}$ $=\frac{f_{uv}}{|f_{v}|}-\frac{f_{v}(|f_{v}|)_{u}}{|f_{v}|^{2}}=\frac{1}{|f_{v}|}\left(\Gamma^{1}_{12}f_{u}+\Gamma^{2}_{12}f_{v}\right)-\Gamma^{2}_{12}\bm{e}_{2}$
$=\frac{E_{v}}{2E|f_{v}|}f_{u}=-\frac{(\kappa_{1})_{v}}{(\kappa_{1}-\kappa_{2})|f_{v}|}f_{u}=-\frac{(\kappa_{1})_{v}|f_{u}|}{(\kappa_{1}-\kappa_{2})|f_{v}|}\bm{e}_{1}.$
By similar calculations, we get
$(\bm{e}_{1})_{v}$ $=\frac{f_{uv}}{|f_{u}|}-\frac{\left\langle{f_{u}},{f_{uv}}\right\rangle}{|f_{u}|^{2}}=\frac{\Gamma^{2}_{12}}{|f_{u}|}f_{v}=\frac{G_{u}}{2G|f_{u}|}f_{v}=\frac{(\kappa_{2})_{u}|f_{v}|}{(\kappa_{1}-\kappa_{2})|f_{u}|}\bm{e}_{2},$
$(\bm{e}_{2})_{v}$ $=\frac{f_{vv}}{|f_{v}|}-\Gamma^{2}_{22}\bm{e}_{2}=\frac{1}{|f_{v}|}\left(\Gamma^{1}_{22}f_{u}+N\nu\right)=-\frac{G_{u}}{2E|f_{v}|}f_{u}+\frac{N}{|f_{v}|}\nu$
$=-\frac{(\kappa_{2})_{u}G}{(\kappa_{1}-\kappa_{2})E|f_{v}|}f_{u}+\kappa_{2}|f_{v}|\nu=-\frac{(\kappa_{2})_{u}|f_{v}|}{(\kappa_{1}-\kappa_{2})|f_{u}|}\bm{e}_{1}+\kappa_{2}|f_{v}|\nu.$
Thus we have the assertions. ∎
By this lemma, it holds that
$(\bm{e}_{1})_{u}\times(\bm{e}_{1})_{v}=-\frac{\kappa_{1}(\kappa_{2})_{u}|f_{v}|}{\kappa_{1}-\kappa_{2}}\bm{e}_{1},\quad(\bm{e}_{2})_{u}\times(\bm{e}_{2})_{v}=-\frac{\kappa_{2}(\kappa_{1})_{v}|f_{u}|}{\kappa_{1}-\kappa_{2}}\bm{e}_{2}.$
Therefore sets of singularities of $\bm{e}_{1}$ and $\bm{e}_{2}$ are
(2.4) $S(\bm{e}_{1})$ $=\{q\in U\mid\kappa_{1}(q)=0\}\cup\{q\in U\mid(\kappa_{2})_{u}(q)=0\},$
$S(\bm{e}_{2})$ $=\{q\in U\mid\kappa_{2}(q)=0\}\cup\{q\in U\mid(\kappa_{1})_{v}(q)=0\},$
respectively. Since $S(\nu)=\kappa_{1}^{-1}(0)\cup\kappa_{2}^{-1}(0)$, if $p\in S(\nu)$, then $p\in S(\bm{e}_{1})$ or $p\in S(\bm{e}_{2})$. See [5] for singularities of frame mappings on regular surfaces.
2.2. Fronts
To investigate parallel surfaces below, we here recall the class of surface with certain singularities called fronts ([1, 28]). Let $f\colon U\to\mathbb{R}^{3}$ be a $C^{\infty}$ map, where $U\subset\mathbb{R}^{2}$ is an open set. Then $f$ is said to be a front if there exists a $C^{\infty}$ map $\nu\colon U\to\mathbb{S}^{2}$ such that
  • $\left\langle{df_{q}(\bm{v})},{\nu(q)}\right\rangle=0$ for any $q\in U$ and $\bm{v}\in T_{q}\mathbb{R}^{2}$, and
  • the pair $(f,\nu)\colon U\to\mathbb{R}^{3}\times\mathbb{S}^{2}$ gives an immersion.
We call the map $\nu$ a unit normal vector field of $f$. By definition, an immersion $f\colon U\to\mathbb{R}^{3}$ is a front.
Let $f\colon U\to\mathbb{R}^{3}$ be a front with a unit normal vector field $\nu$. Then a point $p\in U$ is said to be a singular point of $f$ if $f$ is not an immersion at $p$, that is, $\operatorname{rank}df_{p}<2$. We denote by $S(f)$ the set of singular points of $f$. We define a function $\lambda\colon U\to\mathbb{R}$ by
(2.5) $\lambda(u,v)=\det(f_{u},f_{v},\nu)(u,v).$
We then call $\lambda$ the signed area density function of $f$. It is clear that $S(f)=\lambda^{-1}(0)$ holds. Moreover, a function $\tilde{\lambda}$ which is non-zero proportional to $\lambda$ is called an identifier of singularities of $f$. A singular point $p\in S(f)$ of $f$ is said to be of non-degenerate if $(\tilde{\lambda}_{u},\tilde{\lambda}_{v})(p)\neq(0,0)$. Otherwise, it is said to be of degenerate. For a non-degenerate singular point $p\in S(f)$, there exist a neighborhood $V(\subset U)$ of $p$ and a regular curve $\gamma\colon(-\varepsilon,\varepsilon)\to V$ with $\gamma(0)=p$ such that $\tilde{\lambda}(\gamma(t))=0$ for $t\in(-\varepsilon,\varepsilon)$. This curve $\gamma$ is called a singular curve for $f$ passing through $p=\gamma(0)$. On the other hand, for a singular point $p\in S(f)$ satisfying $\operatorname{rank}df_{p}=1$, there exists a non-zero vector field $\eta$ on $U$ such that $\eta_{p}$ generates the kernel of $df_{p}$. This vector field $\eta$ is called a null vector field for $f$.
Definition 2.4.
Let $f_{i}\colon(U_{i},p_{i})\to\mathbb{R}^{3}$ ($i=1,2$) be $C^{\infty}$ map germs. Then $f_{1}$ and $f_{2}$ are $\cal{A}$-equivalent if there exist diffeomorphism germs $\varphi\colon(U_{1},p_{1})\to(U_{2},p_{2})$ and $\Phi\colon(\mathbb{R}^{3},f_{1}(p_{1}))\to(\mathbb{R}^{3},f_{2}(p_{2}))$ such that $\Phi\circ f_{1}=f_{2}\circ\varphi$.
Definition 2.5.
Let $f\colon U\to\mathbb{R}^{3}$ be a front. Then
  1. (1)
    $f$ is a cuspidal edge at $p(\in U)$ if $f$ is $\cal{A}$-equivalent to the germ $(u,v)\mapsto(u,v^{2},v^{3})$ at the origin;
  2. (2)
    $f$ is a swallowtail at $p$ if $f$ is $\cal{A}$-equivalent to the germ $(u,v)\mapsto(u,4v^{3}+2uv,3v^{4}+uv^{2})$ at the origin;
  3. (3)
    $f$ is a cuspidal butterfly at $p$ if $f$ is $\cal{A}$-equivalent to the germ $(u,v)\mapsto(u,5v^{4}+2uv,4v^{5}+uv^{2})$ at the origin;
  4. (4)
    $f$ is a cuspidal lips at $p$ if $f$ is $\cal{A}$-equivalent to the germ $(u,v)\mapsto(u,2v^{3}+u^{2}v,3v^{4}+u^{2}v^{2})$ at the origin;
  5. (5)
    $f$ is a cuspidal beaks at $p$ if $f$ is $\cal{A}$-equivalent to the germ $(u,v)\mapsto(u,2v^{3}-u^{2}v,3v^{4}-u^{2}v^{2})$ at the origin.
These are rank one singularities of fronts. For these singularities, the following criteria are known (see [13, 11, 12, 23, 28]).
Fact 2.6.
Let $f\colon U\to\mathbb{R}^{3}$ be a front and $p\in U$ a rank one singular point of $f$. Then
  1. (1)
    $f$ is a cuspidal edge at $p$ if and only if $\eta\tilde{\lambda}(p)\neq 0;$
  2. (2)
    $f$ is a swallowtail at $p$ if and only if $d\tilde{\lambda}(p)\neq 0$, $\eta\tilde{\lambda}(p)=0$ and $\eta^{2}\tilde{\lambda}(p)\neq 0;$
  3. (3)
    $f$ is a cuspidal butterfly at $p$ if and only if $d\tilde{\lambda}(p)\neq 0$, $\eta\tilde{\lambda}(p)=\eta^{2}\tilde{\lambda}(p)=0$ and $\eta^{3}\tilde{\lambda}(p)\neq 0;$
  4. (4)
    $f$ is a cuspidal lips at $p$ if and only if $d\tilde{\lambda}(p)=0$ and $\det\operatorname{Hess}(\tilde{\lambda})(p)>0;$
  5. (5)
    $f$ is a cuspidal beaks at $p$ if and only if $d\tilde{\lambda}(p)=0$, $\eta^{2}\tilde{\lambda}(p)\neq 0$ and $\det\operatorname{Hess}(\tilde{\lambda})(p)<0$.
Here $\tilde{\lambda}$ is an identifier of singularities of $f$, $\eta$ a null vector field, $\eta^{k}\tilde{\lambda}$ is the $k$-th order directional derivative of $\tilde{\lambda}$ by $\eta$, and $\det\operatorname{Hess}(\tilde{\lambda})(p)$ is the Hessian of $\tilde{\lambda}$ at $p$.
2.2.1. Limiting normal curvature and singular curvature
Let $f\colon U\to\mathbb{R}^{3}$ be a front and $\nu$ its unit normal vector. Assume that $f$ is a cuspidal edge at $p$. Then one can find a singular curve $\gamma(t)$ passing through $p=\gamma(0)$ and a null vector field $\eta$ around $p$. We set $\hat{\gamma}=f\circ\gamma$. Then along $\gamma$, we can define the following geometric invariants:
(2.6) $\kappa_{\nu}(t)$ $=\frac{\left\langle{\hat{\gamma}^{\prime\prime}(t)},{\nu(\gamma(t))}\right\rangle}{|\hat{\gamma}^{\prime}(t)|^{2}}=-\frac{\left\langle{\hat{\gamma}^{\prime}(t)},{\nu^{\prime}(\gamma(t))}\right\rangle}{|\hat{\gamma}^{\prime}(t)|^{2}},$
$\kappa_{s}(t)$ $=\operatorname{sgn}(\eta\lambda\cdot\det(\gamma^{\prime},\eta))\frac{\det(\hat{\gamma}^{\prime}(t),\hat{\gamma}^{\prime\prime}(t),\nu(\gamma(t)))}{|\hat{\gamma}^{\prime}(t)|^{3}},$
where $\lambda$ is the signed area density function. We call $\kappa_{\nu}$ and $\kappa_{s}$ the limiting normal curvature and the singular curvature, respectively ([16, 22]).
It is known that $\kappa_{\nu}$ relates to boundedness of the Gaussian curvature near a cuspidal edge. In fact, Saji, Umehara and Yamada [22] showed that the Gaussian curvature of $f$ is bounded on a sufficiently small neighborhood of $p$ if and only if $\kappa_{\nu}\equiv 0$ along a singular curvature $\gamma$ passing through $p$ (see Figure 1).
(a) $\kappa_{\nu}=0$
(b) $\kappa_{\nu}\neq 0$
Figure 1. Cuspidal edges with vanishing $\kappa_{\nu}$ (left) and non-vanishing $\kappa_{\nu}$ (right).
On the other hand, $\kappa_{s}$ corresponds to convexity or concavity of a cuspidal edge (see Figure 2).
(a) $\kappa_{s}>0$
(b) $\kappa_{s}<0$
Figure 2. Cuspidal edges with positive $\kappa_{s}$ (left) and negative $\kappa_{s}$ (right).
In particular, for the case where the Gaussian curvature $K$ is bounded near a cuspidal edge $p$, if $K$ takes non-negative values around $p$, then $\kappa_{s}$ takes non-positive values along the singular curve passing through $p$ ([22, Theorem 3.1]).
2.3. Rational order of functions
[…]
Let $h_{1},h_{2}\colon(\mathbb{R}^{2},p)\to\mathbb{R}$ be function germs. Let $Z=h_{2}^{-1}(0)$. Define a functions germ $g=h_{1}/h_{2}\colon(\mathbb{R}^{2}\setminus Z,p)\to\mathbb{R}$. Then the rational order $\operatorname{ord}_{p}{g}$ of $g$ at $p$ is defined by
$\operatorname{ord}_{p}{g}\coloneqq\operatorname{ord}_{p}{h_{1}}-\operatorname{ord}_{p}{h_{2}}.$
[…]
Example 2.7.
Let $h_{1},h_{2}\colon(\mathbb{R}^{2},0)\to\mathbb{R}$ be function germs given by
$h_{1}(u,v)=uv,\quad h_{2}(u,v)=u^{2}-v^{2}.$
[…]
$g(u,v)=\frac{h_{1}(u,v)}{h_{2}(u,v)}=\frac{uv}{u^{2}-v^{2}}.$
This cannot be extended as a bounded function on $(\mathbb{R}^{2},0)$. However, the rational order $\operatorname{ord}_{0}{g}$ of $g$ at 0 is
$\operatorname{ord}_{0}{g}=2-2=0.$
This implies that $g$ is rationally bounded at 0.
3. Parallel surfaces
We consider the parallel surfaces of a regular surface $f\colon U\to\mathbb{R}^{3}$ without umbilic point. For a fixed $t\in\mathbb{R}\setminus\{0\}$, a parallel surface $f^{t}$ of $f$ at distance $t$ is defined as
(3.1) $f^{t}(u,v)\coloneqq f(u,v)+t\nu(u,v),$
where $\nu$ is a unit normal vector to $f$. Let us take curvature line coordinates $(u,v)$. Then we see that
(3.2) $f^{t}_{u}=(1-t\kappa_{1})f_{u},\quad f^{t}_{v}=(1-t\kappa_{2})f_{v}$
by Rodrigues’ formula. Thus $\nu$ can be also a unit normal vector to $f^{t}$. Moreover, $f^{t}$ is a front for any $t\in\mathbb{R}\setminus\{0\}$ ([28]). By (2.1), the first and second fundamental forms of $f^{t}$ are
(3.3) $\operatorname{I}^{t}=(1-t\kappa_{1})^{2}Edu^{2}+(1-t\kappa_{2})^{2}Gdv^{2},\quad\operatorname{II}^{t}=(1-t\kappa_{1})\kappa_{1}Edu^{2}+(1-t\kappa_{2})\kappa_{2}Gdv^{2}.$
Therefore the Gaussian and mean curvatures for $f^{t}$ are
(3.4) $K^{t}=\frac{\kappa_{1}\kappa_{2}}{(1-t\kappa_{1})(1-t\kappa_{2})}=\frac{K}{1-2tH+t^{2}K},\quad H^{t}=\frac{H-tK}{1-2tH+t^{2}K},$
where $K$ and $H$ are the Gaussian and mean curvatures of $f$, respectively.
For each fixed $t\in\mathbb{R}\setminus\{0\}$, the set of singular points $S(f^{t})$ of $f^{t}$ is
$S(f^{t})=\bigcup_{i=1}^{2}\{(u,v)\in U\mid\kappa_{i}(u,v)=1/t\}.$
Since $f$ has no umbilic points,
$\bigcap_{i=1}^{2}\{(u,v)\in U\mid\kappa_{i}(u,v)=1/t\}=\emptyset.$
If we take $t=1/\kappa_{1}(p)$, $S(f^{t})$ coincides with $\kappa_{1}^{-1}(\kappa_{1}(p))$. This implies that the constant principal curvature lines form the set of singular points of a parallel surface (see [8]).
3.1. Singularities of parallel surfaces
In the following of this section, we fix $t=1/\kappa_{1}(p)$. Then we see that $f^{t}_{u}(p)=0$ and $f^{t}_{v}(p)\neq 0$ by (3.2). Thus a point $p$ is a corank one singular point of $f^{t}$. Moreover, we set
(3.5) $\tilde{\lambda}(u,v)=\kappa_{1}(u,v)-\kappa_{1}(p).$
Then $\tilde{\lambda}^{-1}(0)=S(f^{t})$. Thus $\tilde{\lambda}$ is an identifier of singularities of $f^{t}$. In addition, $\partial_{u}$ give a null vector field for $f^{t}$. By Fact 2.6, we have the following immediately.
Proposition 3.1 ([8, Theorems 3.4 and 3.5]).
Let $f\colon U\to\mathbb{R}^{3}$ be a regular surface without umbilic point and $f^{t}$ the parallel surface of $f$ at distance $t=1/\kappa_{1}(p)$. Take a curvature line coordinate system $(u,v)$ on $U$ for $f$.
  1. (1)
    $f^{t}$ is a cuspidal edge at $p$ if and only if $(\kappa_{1})_{u}\neq 0$ at $p$.
  2. (2)
    $f^{t}$ is a swallowtail at $p$ if and only if $(\kappa_{1})_{u}=0$, $(\kappa_{1})_{v}\neq 0$ and $(\kappa_{1})_{uu}\neq 0$ at $p$.
  3. (3)
    $f^{t}$ is a cuspidal butterfly at $p$ if and only if $(\kappa_{1})_{u}=(\kappa_{1})_{uu}=0$, $(\kappa_{1})_{v}\neq 0$ and $(\kappa_{1})_{uuu}\neq 0$ at $p$.
  4. (4)
    $f^{t}$ is a cuspidal lips at $p$ if and only if $(\kappa_{1})_{u}=(\kappa_{1})_{v}=0$ and $\det\operatorname{Hess}(\kappa_{1})>0$ at $p$.
  5. (5)
    $f^{t}$ is a cuspidal beaks at $p$ if and only if $(\kappa_{1})_{u}=(\kappa_{1})_{v}=0$, $(\kappa_{1})_{uu}\neq 0$ and $\det\operatorname{Hess}(\kappa_{1})<0$ at $p$.
Corollary 3.2.
With the same situations in Proposition 3.1, we have the following.
  1. (1)
    If the parallel surface $f^{t}$ is one of a cuspidal edge, a swallowtail or a cuspidal butterfly at $p$, then $\operatorname{ord}_{p}\tilde{\lambda}=1$, where $\tilde{\lambda}$ is the function as in (3.5).
  2. (2)
    If $f^{t}$ is either a cuspidal lips or a cuspidal beaks at $p$, then $\operatorname{ord}_{p}\tilde{\lambda}(p)=2$.
Proof.
We take a curvature line coordinate system $(u,v)$. We assume that $p$ is the origin 0 of $\mathbb{R}^{2}$. Then we remark that
[…]
hold ([10]).
[…]
$\tilde{\lambda}(u,v)=uh_{1}(u,v)+vh_{2}(u,v).$
[…]
3.2. Behavior of the Gaussian curvature
We consider behavior of the Gaussian curvature $K^{t}$ of the parallel surface $f^{t}$ of a regular surface $f$ at a distance $t=1/\kappa_{1}(p)$. When $f^{t}$ has non-degenerate singularity, the following assertion holds.
Proposition 3.3.
Let $f\colon U\to\mathbb{R}^{3}$ be a regular surface and $f^{t}$ the parallel surface with $t=1/\kappa_{1}(p)$. Suppose that $f^{t}$ is one of a cuspidal edge, a swallowtail or a cuspidal butterfly at $p$. Then the Gaussian curvature $K^{t}$ is rationally bounded at $p$ if and only if $\kappa_{2}$ vanishes at $p$. Moreover, $K^{t}$ is bounded near $p$ if and only if $\kappa_{2}$ vanishes on $S(f^{t})$.
Proof.
By [16, Corollary C], $K^{t}$ is rationally bounded at $p$ if and only if the Gauss map $\nu$ of $f^{t}$ is singular at $p$. This is equivalent to the condition that $\kappa_{2}(p)=0$ because $\kappa_{1}(p)\neq 0$. Thus the first assertion holds. Moreover, $K^{t}$ is bounded near $p$ if and only if the second fundamental form of $f^{t}$ vanishes on $S(f^{t})$ ([22, Theorem 3.1]). By (3.3), it holds that
$\operatorname{II}^{t}=(1-t\kappa_{2})\kappa_{2}dv^{2}$
on $S(f^{t})$. Thus the second assertion holds. ∎
Proposition 3.4.
Let $f\colon U\to\mathbb{R}^{3}$ be a regular surface whose parallel surface $f^{t}$ is a cuspidal lips or cuspidal beaks at $p$, where $t=1/\kappa_{1}(p)$. Then the Gaussian curvature $K^{t}$ of $f^{t}$ is rationally bounded at $p$ if and only if $\operatorname{ord}_{p}\kappa_{2}=2$. Moreover, $K^{t}$ is bounded near $p$ if and only if the set of parabolic points $\kappa_{2}^{-1}(0)$ contains $S(f^{t})$.
Proof.
By the assumption and Corollary 3.2, we see that the function $\tilde{\lambda}$ as in (3.5) satisfies $\operatorname{ord}_{p}\tilde{\lambda}=2$. By (3.4) and $t=1/\kappa_{1}(p)$, the Gaussian curvature $K^{t}$ of $f^{t}$ can be written as
$K^{t}=\frac{\kappa_{1}\kappa_{2}}{t^{2}(\kappa_{1}(p)-\kappa_{1})(\kappa_{1}(p)-\kappa_{2})}=\kappa_{1}(p)^{2}\frac{\kappa_{1}\kappa_{2}}{\tilde{\lambda}(\kappa_{2}-\kappa_{1}(p))},$
it follows that the rational order $\operatorname{ord}_{p}K^{t}$ of $K^{t}$ satisfies $\operatorname{ord}_{p}K^{t}=0$ at $p$ if and only if $\operatorname{ord}_{p}\kappa_{2}=2$. This implies that $K^{t}$ is rationally bounded at $p$ if and only if $\operatorname{ord}_{p}\kappa_{2}=2$. The boundedness is also proved. ∎
3.3. The limiting normal curvature
We investigate the limiting normal curvature at a singular point of $f^{t}$.
Proposition 3.5.
Let $f\colon U\to\mathbb{R}^{3}$ be a regular surface without umbilic point and $\nu$ its unit normal vector. Suppose that the principal curvature $\kappa_{1}$ does not vanish on $U$. If the parallel surface $f^{t}$ with $t=1/\kappa_{1}(p)$ satisfies $\operatorname{rank}df^{t}_{p}=1$, then the limiting normal curvature $\kappa_{\nu}^{t}$ of $f^{t}$ satisfies
$\kappa_{\nu}^{t}=\frac{\kappa_{1}\kappa_{2}}{\kappa_{1}-\kappa_{2}}$
at $p$.
Proof.
We take a curvature line coordinates $(u,v)$ on $U$. We first consider the case where the parallel surface $f^{t}$ is a cuspidal edge at $p$. An identifier of singularities $\tilde{\lambda}$ for $f^{t}$ is given by (3.5). By Proposition 3.1, there exists a function $u(v)$ such that $\tilde{\lambda}(u(v),v)=0$. We set a curve $c(v)=(u(v),v)$, and define $\hat{c}(v)=f^{t}(u(v),v)$. Direct calculations show
(3.6) $\hat{c}^{\prime}(v)$ $=f^{t}_{u}(c(v))u^{\prime}(v)+f^{t}_{v}(c(v))=\left(1-\frac{\kappa_{2}(c(v))}{\kappa_{1}(c(v))}\right)f_{v}(c(t)),$
$\hat{\nu}^{\prime}(v)$ $=-\kappa_{1}(c(v))f_{u}(c(v))u^{\prime}(v)-\kappa_{2}(c(v))f_{v}(c(v)),$
where $^{\prime}=d/dv$ and $\hat{\nu}(v)=\nu(u(v),v)$. Thus we have
$\kappa_{\nu}^{t}(v)=-\frac{\left\langle{\hat{c}^{\prime}(v)},{\hat{\nu}^{\prime}(v)}\right\rangle}{|\hat{c}^{\prime}(v)|^{2}}=\frac{\kappa_{1}(c(v))\kappa_{2}(c(v))}{\kappa_{1}(c(v))-\kappa_{2}(c(v))}.$
We next show the case where $p$ is a non-degenerate singular point of $f^{t}$ other than a cuspidal edge. In this case, we get $(\kappa_{1})_{u}(p)=(\kappa_{1})_{uu}(p)=0$ and $(\kappa_{1})_{v}(p)\neq 0$. Thus by the implicit function theorem, there exists a function $v(u)$ such that $\kappa_{1}(u,v(u))-\kappa_{1}(p)=0$. Let us set $c(u)=(u,v(u))$ and $\hat{c}(u)=f^{t}(u,v(u))$. Then it holds that
$\dot{\hat{c}}(u)$ $=f^{t}_{u}(c(u))+f^{t}_{v}(c(u))\dot{v}(u)=\left(1-\frac{\kappa_{2}(c(u))}{\kappa_{1}(c(u))}\right)f_{v}(c(u))\dot{v}(u),$
$\dot{\hat{\nu}}(u)$ $=-\kappa_{1}(c(u))f_{u}(c(u))-\kappa_{2}(c(u))f_{v}(c(u))\dot{v}(u),$
where $\hat{\nu}(u)=\nu(u,v(u))$ and $\dot{}=d/du$. Therefore we get
$|\dot{\hat{c}}(u)|^{2}$ $=\frac{(\kappa_{1}(c(u))-\kappa_{2}(c(u)))^{2}}{\kappa_{1}(c(u))^{2}}\dot{v}(u)^{2}G(c(u)),$
$-\left\langle{\dot{\hat{c}}(u)},{\dot{\hat{\nu}}(u)}\right\rangle$ $=\frac{\kappa_{2}(c(u))(\kappa_{1}(c(u))-\kappa_{2}(c(u)))}{\kappa_{1}(c(u))}\dot{v}(u)^{2}G(c(u)).$
Hence we have
$\kappa_{\nu}^{t}(p)=\lim_{u\to u_{0}}-\frac{\left\langle{\dot{\hat{c}}(u)},{\dot{\hat{\nu}}(u)}\right\rangle}{|\dot{\hat{c}}(u)|^{2}}=\frac{\kappa_{1}(p)\kappa_{2}(p)}{\kappa_{1}(p)-\kappa_{2}(p)}.$
We finally deal with the case where $p$ is a degenerate corank one singular point of $f^{t}$. Since $f^{t}_{u}(p)=0$ and $f^{t}_{v}(p)\neq 0$, the limiting normal curvature can be calculated as
$\kappa_{\nu}^{t}(p)=\frac{\left\langle{f^{t}_{vv}},{\nu(p)}\right\rangle}{|f^{t}_{v}(p)|^{2}}$
(see [16, p.253]). By (3.2), we have
$f^{t}_{vv}=-t(\kappa_{2})_{v}f_{v}+(1-t\kappa_{2})f_{vv}.$
Thus we get
$|f^{t}_{v}(p)|^{2}=\frac{(\kappa_{1}(p)-\kappa_{2}(p))^{2}}{\kappa_{1}(p)^{2}}G(p),\quad\left\langle{f^{t}_{vv}(p)},{\nu(p)}\right\rangle=\frac{\kappa_{2}(p)(\kappa_{1}(p)-\kappa_{2}(p))}{\kappa_{1}(p)}G(p),$
where we used $N(p)=\kappa_{2}(p)G(p)$. Therefore $\kappa_{\nu}^{t}(p)$ can be written by
$\kappa_{\nu}^{t}(p)=\frac{\kappa_{1}(p)\kappa_{2}(p)}{\kappa_{1}(p)-\kappa_{2}(p)}.$
Hence we have the assertions. ∎
By this proposition, we get the following.
Corollary 3.6.
With the same assumptions in Proposition 3.5, the limiting normal curvature $\kappa_{\nu}^{t}$ of $f^{t}$ vanishes at $p$ if and only if $\nu$ is singular at $p$. Moreover, $\bm{e}_{2}$ is also singular at $p$.
Proof.
By Proposition 3.5, $\kappa_{\nu}^{t}(p)=0$ if and only if $\kappa_{2}(p)=0$ under the assumptions. On the other hand, $\nu$ is singular at $p$ if and only if $\kappa_{2}(p)=0$ in this case. Moreover, by (2.4), $\bm{e}_{2}$ is singular at $p$ if $\kappa_{2}(p)=0$. Thus we get the conclusions. ∎
4. Focal surfaces
Let $f\colon U\to\mathbb{R}^{3}$ be a regular surface without umbilic point parametrized by a curvature line coordinate system $(u,v)$. Let $\kappa_{i}$ $(i=1,2)$ be principal curvatures for $f$. Then the maps
(4.1) $C_{i}\coloneqq f+\frac{1}{\kappa_{i}}\nu\colon U\setminus\kappa_{i}^{-1}(0)\to\mathbb{R}^{3}\quad(i=1,2)$
are called focal surfaces (or caustics) of $f$. By direct calculations, we see that
(4.2) $(C_{1})_{u}$ $=-\frac{(\kappa_{1})_{u}}{\kappa_{1}^{2}}\nu,\quad(C_{1})_{v}=\left(1-\frac{\kappa_{2}}{\kappa_{1}}\right)f_{v}-\frac{(\kappa_{1})_{v}}{\kappa_{1}^{2}}\nu,$
$(C_{2})_{u}$ $=\left(1-\frac{\kappa_{1}}{\kappa_{2}}\right)f_{u}-\frac{(\kappa_{2})_{u}}{\kappa_{2}^{2}}\nu,\quad(C_{2})_{v}=-\frac{(\kappa_{2})_{v}}{\kappa_{2}^{2}}\nu.$
Therefore $\bm{e}_{1}$ (resp. $\bm{e}_{2}$) as in (2.3) can be taken as a unit normal vector to $C_{1}$ (resp. $C_{2}$). Thus focal surfaces are frontals in this case. Moreover, the signed area density functions for $C_{1}$ and $C_{2}$ are
(4.3) $\lambda^{C_{1}}$ $=\det((C_{1})_{u},(C_{1})_{v},\bm{e}_{1})=\frac{(\kappa_{1})_{u}}{\kappa_{1}^{2}}\left(1-\frac{\kappa_{2}}{\kappa_{1}}\right)|f_{v}|,$
$\lambda^{C_{2}}$ $=\det((C_{2})_{u},(C_{2})_{v},\bm{e}_{2})=\frac{(\kappa_{2})_{v}}{\kappa_{2}^{2}}\left(1-\frac{\kappa_{1}}{\kappa_{2}}\right)|f_{u}|,$
respectively. Since $\kappa_{1}\neq\kappa_{2}$ on $U$, the sets of singular points of $C_{i}$ ($i=1,2$) are
$S(C_{1})=\{q\in U\mid(\kappa_{1})_{u}(q)=0\},\quad S(C_{2})=\{q\in U\mid(\kappa_{2})_{v}(q)=0\}$
(cf. [10, 20]). For a point $p\in S(C_{1})$ (resp. $p\in S(C_{2})$), we see that $\operatorname{rank}d(C_{1})_{p}=1$ (resp. $\operatorname{rank}d(C_{2})_{p}=1$). Thus one can take $\eta^{C_{1}}=\partial_{u}$ (resp. $\eta^{C_{2}}=\partial_{v}$) as a null vector field for $C_{1}$ (resp. $C_{2}$). Moreover, by Lemma 2.3, it holds that $(\bm{e}_{1})_{u}\neq 0$ (resp. $(\bm{e}_{2})_{v}\neq 0$) at $p\in S(C_{1})$ (resp. $p\in S(C_{2})$). Hence $C_{1}$ (resp. $C_{2}$) is a front at $p\in S(C_{1})$ (resp. $p\in S(C_{2})$) (see [28]). For singularities of focal surfaces, we have the following.
Proposition 4.1 (cf. [26]).
  1. (1)
    When the parallel surface $f^{t}$ with $t=1/\kappa_{i}(p)$ is a cuspidal edge at $p$, then $C_{i}$ is regular at $p$.
  2. (2)
    When $f^{t}$ is either a swallowtail, a cuspidal lips or a cuspidal beaks at $p$, then $C_{i}$ is a cuspidal edge at $p$.
Proof.
We show the case of $C_{1}$. Let us take a curvature line coordinate system $(u,v)$. Assume that the parallel surface $f^{t}$ with $t=1/\kappa_{1}(p)$ of $f$ is a cuspidal edge at $p$. Then by Proposition 3.1, it holds that $(\kappa_{1})_{u}(p)\neq 0$. Thus $p\notin S(C_{1})$. This implies that the first assertion holds. Next, we assume that $f^{t}$ is one of a swallowtail, a cuspidal lips or a cuspidal beaks. Then by Proposition 3.1, it holds that
$(\kappa_{1})_{u}(p)=0,\quad(\kappa_{1})_{uu}(p)\neq 0,\quad(\kappa_{1})_{v}(p)\neq 0.$
On the other hand, since an identifier of singularities for $C_{1}$ can be taken as $\tilde{\lambda}^{C_{1}}=(\kappa_{1})_{u}$ and $\eta^{C_{1}}=\partial_{u}$ is a null vector field for $C_{1}$, we have $\eta^{C_{1}}\tilde{\lambda}^{C_{1}}=(\kappa_{1})_{uu}$, which does not vanish at $p$ by the assumption. Hence $C_{1}$ is a cuspidal edge at $p$ by Fact 2.6. Therefore we get the second assertion.
For the case of $C_{2}$, one can show in a similar way. ∎
Remark 4.2.
This proposition implies that if $f$ is a front with a cuspidal lips or cuspidal beaks, then its (suitable) focal surface is a cuspidal edge at the corresponding point (cf. [26]).
4.1. Geometric invariants at cuspidal edges of focal surfaces
We investigate geometric invariants of focal surfaces at cuspidal edges. Let $f\colon U\to\mathbb{R}^{3}$ be a regular surface parametrized by a curvature line coordinate system $(u,v)$. We then assume that $\kappa_{2}(p)=0$ and the focal surface $C_{1}$ is a cuspidal edge at $p\in U$. In this case, we shall consider geometric invariants called the limiting normal curvature $\kappa_{\nu}^{C_{1}}$ and the singular curvature $\kappa_{s}^{C_{1}}$ of $C_{1}$ at $p$.
By the assumption, $S(C_{1})=((\kappa_{1})_{u})^{-1}(0)$ and $(\kappa_{1})_{uu}(p)\neq 0$ hold. Thus there exists a regular curve $\gamma(v)=(u(v),v)$ ($|v-v_{0}|<\varepsilon$) with $\gamma(v_{0})=p$ such that $(\kappa_{1})_{u}(u(v),v)=0$ by the implicit function theorem. The curve $\gamma(v)$ is a singular curve for $C_{1}$. Thus the curve
(4.4) $\hat{\gamma}(v)=C_{1}(\gamma(v))=C_{1}(u(v),v)$
gives the singular locus of $C_{1}$ near $p$. We investigate the first and second differentials of $\hat{\gamma}$. By (4.2) and $(\kappa_{1})_{u}(u(v),v)=0$, it holds that
(4.5) $\hat{\gamma}^{\prime}(v)$ $=(C_{1})_{v}(\gamma(v))=\left(1-\frac{\kappa_{2}(\gamma(v))}{\kappa_{1}(\gamma(v))}\right)f_{v}(\gamma(v))-\frac{(\kappa_{1})_{v}(\gamma(v))}{\kappa_{1}(\gamma(v))^{2}}\nu(\gamma(v)),$
$\hat{\gamma}^{\prime\prime}(v)$ $=(C_{1})_{uv}(\gamma(v))u^{\prime}(v)+(C_{1})_{vv}(\gamma(v)),$
where $^{\prime}=d/dv$. By (2.1) and (4.2), we have
(4.6) $(C_{1})_{uv}$ $=\left(\frac{2(\kappa_{1})_{u}(\kappa_{1})_{v}}{\kappa_{1}^{3}}-\frac{(\kappa_{1})_{uv}}{\kappa_{1}^{2}}\right)\nu+\frac{\kappa_{2}(\kappa_{1})_{u}}{\kappa_{1}^{2}}f_{v}=-\frac{(\kappa_{1})_{uv}}{\kappa_{1}^{2}}\nu,$
$(C_{1})_{vv}$ $=\left(1-\frac{\kappa_{2}}{\kappa_{1}}\right)f_{vv}+\left(\frac{2\kappa_{2}(\kappa_{1})_{v}}{\kappa_{1}^{2}}-\frac{(\kappa_{2})_{v}}{\kappa_{1}}\right)f_{v}+\left(\frac{2(\kappa_{1})_{v}^{2}}{\kappa_{1}^{3}}-\frac{(\kappa_{1})_{vv}}{\kappa_{1}^{2}}\right)\nu$
$=\left(1-\frac{\kappa_{2}}{\kappa_{1}}\right)\Gamma^{1}_{22}f_{u}+\left(\frac{2\kappa_{2}(\kappa_{1})_{v}}{\kappa_{1}^{2}}-\frac{(\kappa_{2})_{v}}{\kappa_{1}}+\left(1-\frac{\kappa_{2}}{\kappa_{1}}\right)\Gamma^{2}_{22}\right)f_{v}$
$\hskip 142.26378pt+\left(\frac{2(\kappa_{1})_{v}^{2}}{\kappa_{1}^{3}}-\frac{(\kappa_{1})_{vv}}{\kappa_{1}^{2}}+\left(1-\frac{\kappa_{2}}{\kappa_{1}}\right)N\right)\nu$
along $\gamma$. On the other hand, differentiating $(\kappa_{1})_{u}(u(v),v)=0$ by $v$, we get
$(\kappa_{1})_{uu}(\gamma(v))u^{\prime}(v)+(\kappa_{1})_{uv}(\gamma(v))=0.$
Thus we see that
(4.7) $u^{\prime}(v)=-\frac{(\kappa_{1})_{uv}(\gamma(v))}{(\kappa_{1})_{uu}(\gamma(v))}$
near $p$. By (4.5), (4.6) and (4.7), the second derivative of $\hat{\gamma}$ is
(4.8) $\hat{\gamma}^{\prime\prime}(v)=\left(1-\frac{\kappa_{2}(\gamma(v))}{\kappa_{1}(\gamma(v))}\right)\Gamma^{1}_{22}(\gamma(v))f_{u}(\gamma(v))+\left(\frac{2\kappa_{2}(\gamma(v))(\kappa_{1})_{v}(\gamma(v))}{\kappa_{1}(\gamma(v))^{2}}-\frac{(\kappa_{2})_{v}(\gamma(v))}{\kappa_{1}(\gamma(v))}+\left(1-\frac{\kappa_{2}(\gamma(v))}{\kappa_{1}(\gamma(v))}\right)\Gamma^{2}_{22}(\gamma(v))\right)f_{v}(\gamma(v))+\left(-\frac{\det\operatorname{Hess}(\kappa_{1})(\gamma(v))}{\kappa_{1}(\gamma(v))^{2}(\kappa_{1})_{uu}(\gamma(v))}+\frac{2(\kappa_{1})_{v}(\gamma(v))^{2}}{\kappa_{1}(\gamma(v))^{3}}+\left(1-\frac{\kappa_{2}(\gamma(v))}{\kappa_{1}(\gamma(v))}\right)N(\gamma(v))\right)\nu(\gamma(v)),$
where $\det\operatorname{Hess}(\kappa_{1})(\gamma(v))=(\kappa_{1})_{uu}(\gamma(v))(\kappa_{1})_{vv}(\gamma(v))-(\kappa_{1})_{uv}(\gamma(v))^{2}$.
For the limiting normal curvature, we get the following assertion.
Proposition 4.3.
Let $f\colon U\to\mathbb{R}^{3}$ be a regular surface parametrized by a curvature line coordinates $(u,v)$. Suppose that the focal surface $C_{1}$ of $f$ is a cuspidal edge at $p$ and $\kappa_{2}(p)=0$. Then the limiting normal curvature $\kappa_{\nu}^{C_{1}}$ at $p$ is calculated as
$\kappa_{\nu}^{C_{1}}(p)=-\frac{\kappa_{1}(p)^{3}(\kappa_{2})_{u}(p)G(p)}{\sqrt{E(p)}\left(\kappa_{1}(p)^{4}G(p)+(\kappa_{1})_{v}(p)^{2}\right)}$
Proof.
Since $\bm{e}_{1}$ is a unit normal vector to $C_{1}$, the limiting normal curvature of $C_{1}$ at $p$ is
$\kappa_{\nu}^{C_{1}}(p)=\frac{\left\langle{\hat{\gamma}^{\prime\prime}(v_{0})},{\bm{e}_{1}(p)}\right\rangle}{|\hat{\gamma}^{\prime}(v_{0})|^{2}}$
by (2.6), where $\gamma(v)$ is a singular curve for $C_{1}$ with $\gamma(v_{0})=p$ and $\hat{\gamma}=C_{1}\circ\gamma$. By (2.1), (4.5), (4.8) and Lemma 2.1, we have
$|\hat{\gamma}^{\prime}(v_{0})|^{2}$ $=\frac{1}{\kappa_{1}(p)^{4}}\left(\kappa_{1}(p)^{4}G(p)+(\kappa_{1})_{v}(p)^{2}\right),$
$\left\langle{\hat{\gamma}^{\prime\prime}(v_{0})},{\bm{e}_{1}(p)}\right\rangle$ $=\Gamma^{1}_{22}(p)\sqrt{E(p)}=-\frac{G_{u}}{2\sqrt{E(p)}}=-\frac{(\kappa_{2})_{u}(p)G(p)}{\kappa_{1}(p)\sqrt{E(p)}}.$
Thus we obtain the assertion. ∎
Corollary 4.4.
With the same situations in Proposition 4.3, the point $p$ is a $\partial_{u}$-subparabolic point for $f$ if and only if $\kappa_{\nu}^{C_{1}}$ vanishes at $p$. Moreover, $\bm{e}_{1}$ is singular at $p$ when $\kappa_{\nu}^{C_{1}}(p)=0$. Furthermore, the Gaussian curvature $K^{C_{1}}$ of $C_{1}$ is rationally bounded at $p$ in that case.
Proof.
The first assertion follows from Proposition 4.3. We get the second statement by Lemma 2.3. The last assertion holds by (A.6) and [16, Corollary C]. ∎
For the singular curvature $\kappa_{s}^{C_{1}}$ of $C_{1}$, we have the following.
Proposition 4.5.
Let $f\colon U\to\mathbb{R}^{3}$ be a regular surface parametrized by curvature line coordinates $(u,v)$. Suppose that $\kappa_{2}(p)=0$ and the focal surface $C_{1}$ is a cuspidal edge at $p$. Then the singular curvature $\kappa_{s}^{C_{1}}$ of $C_{1}$ is
$\kappa_{s}^{C_{1}}=\operatorname{sgn}((\kappa_{1})_{uu})\frac{\kappa_{1}^{3}\sqrt{G}}{(\kappa_{1})_{uu}\delta^{3}}\left(\kappa_{1}\det\operatorname{Hess}(\kappa_{1})-(\kappa_{1})_{v}(\kappa_{1})_{uu}\left(\Gamma^{2}_{22}\kappa_{1}+2(\kappa_{1})_{v}-(\kappa_{2})_{v}\right)\right)$
at $p$, where we set
$\delta=\sqrt{\kappa_{1}(p)^{4}G(p)+(\kappa_{1})_{v}(p)^{2}}.$
Proof.
Since $\bm{e}_{1}$ is a unit normal vector to $C_{1}$, the singular curvature $\kappa_{s}^{C_{1}}$ at $p$ is
$\kappa_{s}^{C_{1}}(p)=\operatorname{sgn}(\det(\gamma^{\prime},\eta)(v_{0})\cdot\eta^{C_{1}}\lambda^{C_{1}}(p))\frac{\det(\hat{\gamma}^{\prime}(v_{0}),\hat{\gamma}^{\prime\prime}(v_{0}),\bm{e}_{1}(p))}{|\hat{\gamma}^{\prime}(v_{0})|^{3}}.$
By the scalar triple product, it holds that
$\det(\hat{\gamma}^{\prime}(v_{0}),\hat{\gamma}^{\prime\prime}(v_{0}),\bm{e}_{1}(p))=\left\langle{\bm{e}_{1}(p)\times\hat{\gamma}^{\prime}(v_{0})},{\hat{\gamma}^{\prime\prime}(v_{0})}\right\rangle.$
By (4.5), we see that
$\bm{e}_{1}(p)\times\hat{\gamma}^{\prime}(v_{0})=\frac{(\kappa_{1})_{v}(p)}{\kappa_{1}(p)^{2}}\bm{e}_{2}(p)+|f_{v}(p)|\nu(p).$
Thus by (4.8), we have
$\det(\hat{\gamma}^{\prime}(v_{0}),\hat{\gamma}^{\prime\prime}(v_{0}),\bm{e}_{1}(p))$
$=\frac{-|f_{v}(p)|}{\kappa_{1}(p)^{3}(\kappa_{1})_{uu}(p)}\Big{(}\kappa_{1}(p)\det\operatorname{Hess}(\kappa_{1})(p)$
$\qquad\qquad-(\kappa_{1})_{v}(p)(\kappa_{1})_{uu}(p)\left(\kappa_{1}(p)\Gamma^{2}_{22}(p)+2(\kappa_{1})_{v}(p)-(\kappa_{2})_{v}(p)\right)\Big{)}.$
On the other hand, since $\eta^{C_{1}}=\partial_{u}$, we get
$\eta^{C_{1}}\lambda^{C_{1}}(p)=\frac{(\kappa_{1})_{uu}(p)}{\kappa_{1}(p)^{2}}|f_{v}(p)|,\quad\det(\gamma^{\prime},\eta)(v_{0})=-1$
by (4.3). Hence it follows that
$\operatorname{sgn}(\det(\gamma^{\prime},\eta)(v_{0})\cdot\eta^{C_{1}}\lambda^{C_{1}}(p))=-\operatorname{sgn}((\kappa_{1})_{uu}(p)).$
Since
$|\hat{\gamma}^{\prime}(v_{0})|=\frac{1}{\kappa_{1}(p)^{2}}\sqrt{(\kappa_{1})(p)^{4}G(p)+(\kappa_{1})_{v}(p)^{2}}=\frac{\delta}{\kappa_{1}(p)^{2}},$
we have the conclusion. ∎
By the expression as in Proposition 4.5, when the parallel surface $f^{t}$ with $t=1/\kappa_{1}(p)$ is a swallowtail at $p$, $\kappa_{s}^{C_{1}}$ takes either a positive or negative value at $p$. However, when $f^{t}$ is either a cuspidal lips or a cuspidal beaks at $p$, one can notice the following.
Theorem 4.6.
Let $f\colon U\to\mathbb{R}^{3}$ be a regular surface whose parallel surface $f^{t}$ is a cuspidal lips (resp. a cuspidal beaks) at $p\in U$, where $t=1/\kappa_{1}(p)$. Suppose that $\kappa_{2}(p)=0$. Then the singular curvature $\kappa_{s}^{C_{1}}$ at a cuspidal edge $p$ of $C_{1}$ is positive (resp. negative).
Proof.
Let us take a curvature line coordinate system $(u,v)$. Then by Proposition 3.1, it holds that $(\kappa_{1})_{u}(p)=(\kappa_{1})_{v}(p)=0$ and $\det\operatorname{Hess}(\kappa_{1})(p)>0$ (resp. $\det\operatorname{Hess}(\kappa_{1})(p)<0$) when $f^{t}$ is a cuspidal lips (resp. a cuspidal beaks) at $p$. On the other hand, by Proposition 4.5, the singular curvature $\kappa_{s}^{C_{1}}(p)$ is given by
$\kappa_{s}^{C_{1}}(p)=\operatorname{sgn}((\kappa_{1})_{uu}(p))\frac{\det\operatorname{Hess}(\kappa_{1})(p)}{(\kappa_{1})_{uu}(p)\kappa_{1}(p)^{2}G(p)}.$
Therefore $\operatorname{sgn}(\kappa_{s}^{C_{1}}(p))=\operatorname{sgn}(\det\operatorname{Hess}(\kappa_{1})(p))$ holds. Hence we get the assertion. ∎
If the Gaussian curvature $K^{t}$ of $f^{t}$ with $t=1/\kappa_{1}(p)$ is at least ratinally bounded at $p$, then $\kappa_{2}(p)=0$ by Proposition 3.4. Thus we get the following.
Corollary 4.7.
Let $f\colon U\to\mathbb{R}^{3}$ be a regular surface whose parallel surface $f^{t}$ is a cuspidal lips (resp. cuspidal beaks) at $p\in U$, where $t=1/\kappa_{1}(p)$.
  1. (1)
    If the Gaussian curvature $K^{t}$ of $f^{t}$ is rationally bounded at $p$, then the singular curvature $\kappa_{s}^{C_{1}}$ for the focal surface $C_{1}$ of $f$ is positive (resp. negative) at $p$.
  2. (2)
    If the Gaussian curvature $K^{t}$ of $f^{t}$ is bounded near $p$, then $\kappa_{s}^{C_{1}}$ for $C_{1}$ is positive (resp. negative) at $p$.
Example 4.8.
Let $f\colon\mathbb{R}^{2}\to\mathbb{R}^{3}$ be a $C^{\infty}$ map given by
$f(u,v)=\left(u,v,\frac{1}{2}u^{2}+u^{4}+u^{3}v\right).$
Then principal curvatures satisfy
$\kappa_{1}(0,0)=1,\quad\kappa_{2}(0,0)=0.$
Moreover, we see that the parallel surface $f^{1}=f+\nu$ is a cuspidal beaks at $(0,0)$ (see Figure 3). We denote the coefficients of the first and second fundamental forms by $E,F,G,L,M$ and $N$. Note that $N=0$. Then one can take a principal vector $\bm{v}_{1}$ with respect to $\kappa_{1}$ is
$\bm{v}_{1}=(\kappa_{1}G,M-\kappa_{1}F).$
This is a null vector of $C_{1}$. It is known that $(\bm{v}_{1}\kappa_{1})^{-1}(0)$ corresponds to the set of singular points of $C_{1}$ ([20]). We remark that $((\bm{v}_{1}\kappa_{1})_{u},(\bm{v}_{1}\kappa_{1})_{v})=(21,6)\neq(0,0)$ at $(0,0)$. Thus a vector field
$\xi=-(\bm{v}_{1}\kappa_{1})_{v}\partial_{u}+(\bm{v}_{1}\kappa_{1})_{u}\partial_{v}$
is tangent to $(\bm{v}_{1}\kappa_{1})^{-1}(0)$. Then it follows that $\det(\xi{C_{1}},\xi\xi{C_{1}},\bm{e}_{1})(0,0)=15876>0$, where $\bm{e}_{1}=\bm{v}_{1}f/|\bm{v}_{1}f|$ is a unit normal vector to $C_{1}$. On the other hand, it holds that
$\bm{v}_{1}\lambda^{C_{1}}=21\quad\det(\xi,\bm{v}_{1})=-21$
at $(0,0)$. Thus we have
$\kappa_{s}^{C_{1}}(0)=\operatorname{sgn}(\bm{v}_{1}\lambda^{C_{1}}(0,0)\cdot\det(\xi,\bm{v}_{1})(0,0))\frac{\det(\xi{C_{1}},\xi\xi{C_{1}},\bm{e}_{1})}{|\xi{C_{1}}|^{3}}(0,0)=-\frac{12}{7}<0.$
Figure 3. The images of $f$ (blue), $f^{1}$ (purple) and $C_{1}$ (gray).
Example 4.9.
Let $f\colon\mathbb{R}^{2}\to\mathbb{R}^{3}$ be a regular surface given by
$f(u,v)=\left(u,v,\frac{1}{2}u+uv^{2}+u^{4}\right).$
Then
$\nu(u,v)=\frac{(-u-v^{2}-4u^{3},-2uv,1)}{\sqrt{1+4u^{2}v^{2}+(u+v^{2}+4u^{3})^{2}}}$
is a unit normal vector to $f$. By direct calculation, we see that $f^{1}=f+\nu$ is a cuspidal lips at $(0,0)$. Moreover, following similar calculations as in the previous example, it can be observed that the singular curvature $\kappa_{s}^{C_{1}}$ of the focal surface $C_{1}$ is positive at $(0,0)$ (see Figure 4). Indeed, we have $\kappa_{s}^{C_{1}}(0)=8>0$.
Figure 4. The images of $f$ (red), $f^{1}$ (green) and $C_{1}$ (purple).
Appendix A Gaussian curvature of the focal surface
For a surface $f\colon U\to\mathbb{R}^{3}$ parametrized by curvature line coordinates $(u,v)$, we recall the formula for the Gaussian curvature of the focal surface $C_{1}$. By (4.2), the first fundamental form of $C_{1}$ is
(A.1) $\operatorname{I}^{C_{1}}=\frac{(\kappa_{1})_{u}^{2}}{\kappa_{1}^{4}}du^{2}+\frac{2(\kappa_{1})_{u}(\kappa_{1})_{v}}{\kappa_{1}^{4}}dudv+\left(\left(1-\frac{\kappa_{2}}{\kappa_{1}}\right)^{2}G+\frac{(\kappa_{1})_{v}^{2}}{\kappa_{1}^{4}}\right)dv^{2}.$
Thus we get
(A.2) $\det\operatorname{I}^{C_{1}}=\frac{(\kappa_{1})_{u}^{2}}{\kappa_{1}^{4}}\left(1-\frac{\kappa_{2}}{\kappa_{1}}\right)^{2}G=\frac{(\kappa_{1})_{u}^{2}(\kappa_{1}-\kappa_{2})^{2}G}{\kappa_{1}^{6}}.$
On the other hand, the second fundamental form $\operatorname{II}^{C_{1}}$ of $C_{1}$ with respect to $\bm{n}$ is given by
(A.3) $\operatorname{II}^{C_{1}}=-\frac{(\kappa_{1})_{u}L}{\kappa_{1}^{2}|f_{u}|}du^{2}+\frac{G_{u}}{2|f_{u}|}\left(1-\frac{\kappa_{2}}{\kappa_{1}}\right)dv^{2}.$
By Lemma 2.1, we see that
$\frac{G_{u}}{2}=\frac{(\kappa_{2})_{u}G}{\kappa_{1}-\kappa_{2}}.$
Substituting this for (A.3), we obtain
(A.4) $\operatorname{II}^{C_{1}}=-\frac{(\kappa_{1})_{u}L}{\kappa_{1}^{2}|f_{u}|}du^{2}+\frac{(\kappa_{2})_{u}G}{\kappa_{1}|f_{u}|}dv^{2}.$
Thus we have
(A.5) $\det\operatorname{II}^{C_{1}}=-\frac{(\kappa_{1})_{u}(\kappa_{2})_{u}GL}{E\kappa_{1}^{3}}=-\frac{(\kappa_{1})_{u}(\kappa_{2})_{u}G}{\kappa_{1}^{2}}.$
By (A.2) and (A.5), the Gaussian curvature $K^{C_{1}}$ of $C_{1}$ can be calculated as
(A.6) $K^{C_{1}}=\frac{\det\operatorname{II}^{C_{1}}}{\det\operatorname{I}^{C_{1}}}=-\frac{\kappa_{1}^{4}(\kappa_{2})_{u}}{(\kappa_{1})_{u}(\kappa_{1}-\kappa_{2})^{2}}.$
Thus the Gaussian curvature $K^{C_{1}}$ of $C_{1}$ is bounded near a singular point $p$ if and only if $((\kappa_{1})_{u})^{-1}(0)\subset((\kappa_{2})_{u})^{-1}(0)$. In particular, if $K^{C_{1}}$ is bounded and takes a non-zero value at $p$, then $((\kappa_{1})_{u})^{-1}(0)=((\kappa_{2})_{u})^{-1}(0)$.
Assume that the Gaussian curvature of $f$ is a non-zero constant $c$. Then it holds that $\kappa_{2}=c/\kappa_{1}$. Thus we see that $((\kappa_{2})_{u})^{-1}(0)=((\kappa_{1})_{u})^{-1}(0)$. This implies that $K^{C_{1}}$ is bounded near the set of singular points of $C_{1}$. Moreover, by (A.6), the Gaussian curvature $K^{C_{1}}$ of $C_{1}$ is given by
(A.7) $K^{C_{1}}=\frac{c\kappa_{1}^{4}}{(c-\kappa_{1}^{2})^{2}}.$
Therefore if the surface $f$ is a constant positive (resp. negative) Gaussian curvature, then the Gaussian curvature $K^{C_{1}}$ of the caustic takes positive (resp. negative) values.

### v2

@@
 Key words and phrases:
-parallel surface, focal surface, singularity, cuspidal lips, cuspidal beaks, Gaussian curvature
+parallel surface, focal surface, cuspidal edge
 2020 Mathematics Subject Classification:
@@
 1. Introduction
-For a regular surface in Euclidean $3$-space $\mathbb{R}^{3}$, its parallels and focal surfaces can be examined.
-While the initial surface is non-singular, its parallels and focal surfaces generally have singularities ([10]).
-From the viewpoint of singularity theory of smooth mappings, singularities that appear on parallel surfaces are Legendrian singularities, while those on focal surfaces are Lagrangian singularities ([1, 2, 10]).
-In both cases, a cuspidal edge is known to be a generic singularity ([1]).
-It has been established that there is a relationship between singularities of parallels and focal surfaces of a regular surface ([8, 26]).
-See Table 1 for example.
-We remark that parallel surfaces are fronts ([1, 2, 10, 28]).
+Let $f\colon U\to\mathbb{R}^{3}$ be a $C^{\infty}$ regular surface, where $U$ is an open set in $\mathbb{R}^{2}$.
+Then for a fixed $t\in\mathbb{R}$, a parallel surface $f^{t}\colon U\to\mathbb{R}^{3}$ of $f$ is defined by
+$f^{t}(u,v)=f(u,v)+t\nu(u,v)\quad((u,v)\in U),$
+where $\nu$ is the unit normal vector field along $f$ given by $\nu=(f_{u}\times f_{v})/|f_{u}\times f_{v}|$.
+A parallel surface is a (wave) front in $\mathbb{R}^{3}$ ([1, 2, 25]).
+Let $\kappa_{i}$ ($i=1,2$) be principal curvatures of $f$.
+Then $f^{t}$ is singular at $p\in U$ when $t=1/\kappa_{i}(p)$ ($i=1$ or $2$).
+Singularities of parallel surfaces are Legendrian singularities.
+If $p$ is not an umbilic point of $f$, that is, $\kappa_{1}(p)\neq\kappa_{2}(p)$, then $f^{t}$ may be a cuspidal edge, a swallowtail, a cuspidal lips, a cuspidal beaks and a cuspidal butterfly at $p$ (see Definition 2.3; cf.
+[1, Chapters 21 and 22]).
+Characterizations for these singularities are given by Fukui and Hasegawa [5, Theorems 3.4 and 3.5] in terms of geometric properties of the initial surface.
+On the other hand, one can define focal surfaces $C_{i}\colon U\setminus\kappa_{i}^{-1}(0)\to\mathbb{R}^{3}$ ($i=1,2$) associated with $\kappa_{i}$ by
+$C_{i}(u,v)=f(u,v)+\frac{1}{\kappa_{i}(u,v)}\nu(u,v).$
+The focal surface $C_{i}$ corresponds to the locus of centers of principal curvature spheres with respect to $\kappa_{i}$.
+The relationship between the Gaussian curvature of the focal surface and geometric properties of the initial surface is well understood (see [4, page 183]).
+Moreover, it is known that focal surfaces have Lagrangian singularities ([1, Chapter 21]).
+Furthermore, it can be seen that the singularities of parallel surfaces are located on a focal surface.
+The relationships between several singularities of parallel surfaces and of focal surfaces are shown in Table 1 (see Proposition 4.2).
+Table 1. Relation between singularities.
 singularity of parallel surface singularity of focal surface
@@
 cuspidal beaks cuspidal edge
-Table 1. Relation between singularities (see Proposition 4.1).
-Conversely, there are numerous studies on surface with cuspidal edge singularities from the perspective of differential geometry (cf.
-[7, 16, 14, 19, 22, 24, 25]).
+On the other hand, there are various studies on surface with cuspidal edge singularities from differential geometric viewpoint (cf.
+[6, 12, 13, 14, 16, 20, 21, 22, 24]).
 At a cuspidal edge, several geometric invariants are introduced ([12, 13, 20]).
-In particular, geometric invariants called the limiting normal curvature and the singular curvature play a crucial role.
-It has been established that the limiting normal curvature is associated with the boundedness of the Gaussian curvature at a cuspidal edge ([22, 16]).
-While the singular curvature relates to convexity or concavity of cuspidal edge.
-More precisely, the following fact is known ([22, Theorem 3.1]): The Gaussian curvature is bounded near a cuspidal edge if and only if the limiting normal curvature vanishes along the edge.
-Moreover, if the Gaussian curvature takes non-negative values near a cuspidal edge, then the singular curvature is non-positive at the cuspidal edge.
-For a pseudo-spherical front, that is, a surface with certain singularities whose Gaussian curvature is negative constant on the set of regular points, the Gaussian curvature of its focal surface is bounded (cf.
-(A.7)).
-In addition, we see the following assertions ([27]): If the initial pseudo-spherical front has a cuspidal lips (resp.
-cuspidal beaks), then the singular curvature of the focal surface is positive (resp.
-negative) at the corresponding cuspidal edge.
-Thus it is natural to ask that how singularities of fronts (parallel surfaces) relate to geometric properties of cuspidal edges on focal surfaces.
-In this paper, we investigate the limiting normal curvature and the singular curvature at cuspidal edges of focal surfaces of regular surfaces.
-In particular, we deal with a surface whose Gauss map has singularity at some point.
-In that case, the Gaussian curvature of parallel surfaces satisfies certain bounded condition (see Propositions 3.3 and 3.4).
-Moreover, the limiting normal curvature defined at a corank one singular point on the parallel surface vanishes in the case (Proposition 3.5).
-Comparing conditions for singularities of parallel surfaces and focal surfaces, it holds that the focal surface is a cuspidal edge if the parallel surface is a swallowtail, a cuspidal edge or a cuspidal beaks (Proposition 4.1).
-For a curvature line parametrized surfaces, we give explicit representations for the limiting normal curvature and the singular curvature at a cuspidal edge of the focal surface (Propositions 4.3 and 4.5).
-In particular for the singular curvature, we shall show that relation between the sign of the singular curvature at a cuspidal edge of the focal surface and singularities of parallel surface (Theorem 4.6).
+The limiting normal curvature $\kappa_{\nu}$ and singular curvature $\kappa_{s}$ are known as principal geometric invariants at a cuspidal edge.
+The limiting normal curvature is related to the behavior of the Gaussian curvature near a cuspidal edge ([12, Theorem A]; see Figure 1).
+(We remark that the limiting normal curvature can be defied at a rank one singular point of a front ([12, pages 253–254]).)
+(a) $\kappa_{\nu}=0$
+(b) $\kappa_{\nu}\neq 0$
+Figure 1. Cuspidal edges with vanishing $\kappa_{\nu}$ (left) and non-vanishing $\kappa_{\nu}$ (right). Arrows indicate unit normal vectors.
+Conversely, the sign of the singular curvature reflects the concavity or convexity of the cuspidal edge (see Figure 2; [20, Corollary 1.18]).
+(a) $\kappa_{s}>0$
+(b) $\kappa_{s}<0$
+Figure 2. Cuspidal edges with positive $\kappa_{s}$ (left) and negative $\kappa_{s}$ (right).
+Therefore examining the singular curvature at a cuspidal edge on focal surfaces determines its concavity or convexity.
+Consequently, by investigating the properties of the initial surface or parallel surface that determines the sign of the singular curvature at cuspidal edges appearing on the focal surface, it is expected that the geometric properties of the focal surface will become clearer.
+In this paper, we study geometric properties of parallel surfaces and focal surfaces at rank one singular points.
+In particular, we focus on relationships between geometric perspective and behavior of the limiting normal and singular curvatures.
+In Section 3, we investigate parallel surfaces.
+When the initial surface does not have umbilic points, its parallel surfaces may have rank one singularities.
+Thus one can consider the limiting normal curvature at that point on a parallel surface.
+In particular, the following results hold.
+Proposition 1.1.
+Let $f\colon U\to\mathbb{R}^{3}$ be a regular surface without umbilic point and $\nu$ its unit normal vector.
+Suppose that the principal curvature $\kappa_{1}$ (resp.
+$\kappa_{2}$) does not vanish on $U$.
+If the parallel surface $f^{t}$ with $t=1/\kappa_{1}(p)$ (resp.
+$t=1/\kappa_{2}(p)$) satisfies $\operatorname{rank}df^{t}_{p}=1$, then the limiting normal curvature $\kappa_{\nu}^{t}$ of $f^{t}$ satisfies
+$\kappa_{\nu}^{t}=\frac{\kappa_{1}\kappa_{2}}{\kappa_{1}-\kappa_{2}}\quad\left(\textit{resp.}~\kappa_{\nu}^{t}=\frac{\kappa_{1}\kappa_{2}}{\kappa_{2}-\kappa_{1}}\right)$
+at $p$.
+In Section 4, we deal with focal surfaces.
+Especially, we study the limiting normal curvature and the singular curvature at a cuspidal edge on the focal surface.
+The curvature line coordinate system is employed to provide explicit representations of these curvatures.
+As a consequence, we establish a geometric criterion that determines the sign of the singular curvature of the focal surface in terms of the type of singularity of the parallel surface.
+More precisely, we shall show the following.
+Theorem 1.2.
+Let $f\colon U\to\mathbb{R}^{3}$ be a regular surface whose parallel surface $f^{t}$ is a cuspidal lips (resp.
+a cuspidal beaks) at $p\in U$, where $t=1/\kappa_{i}(p)$ ($i=1$ or $2$).
+Suppose that the limiting normal curvature $\kappa_{\nu}^{t}$ of $f^{t}$ vanishes at $p$.
+Then the singular curvature $\kappa_{s}^{C_{i}}$ at a cuspidal edge $p$ of $C_{i}$ is positive (resp.
+negative).
+We note that if the assumption $\kappa_{2}(p)=0$ is removed, the conclusion of Theorem 1.2 does not hold in general (Example 4.7).
 Acknowledgments
 The author thanks Kentaro Saji and Runa Shimada for fruitful discussions and valuable comments.
+The author is also grateful to the referees for careful reading and constructive comments, which improved the manuscript.
 2. Preliminaries
 2.1. Surfaces in Euclidean 3-space
-We here review some fundamental concepts from differential geometry of immersed surfaces in Euclidean $3$-space $\mathbb{R}^{3}$. For details, see [6, 9, 21].
-2.1.1. Surfaces parametrized by curvature line coordinates
+We here review some fundamental concepts from differential geometry of immersed surfaces in Euclidean $3$-space $\mathbb{R}^{3}$.
+For details, see [3, 7, 18].
 Let $U\subset\mathbb{R}^{2}$ be a domain.
-Let $f\colon U\to\mathbb{R}^{3}$ be an immersion and $\nu$ its unit normal vector.
-Then we call the image of $f$ or $f$ itself a surface.
-We assume that $f$ is a surface without umbilic point in the following.
-Then we can take a curvature line coordinate system $(u,v)$ on $U$.
-In this case, it holds that
-$\left\langle{f_{u}},{f_{v}}\right\rangle=\left\langle{\nu_{u}},{f_{v}}\right\rangle=\left\langle{\nu_{v}},{f_{u}}\right\rangle=0,$
-where $\left\langle{\cdot},{\cdot}\right\rangle$ is the canonical inner product on $\mathbb{R}^{3}$, $(\cdot)_{u}=\partial(\cdot)/\partial u$ and $(\cdot)_{v}=\partial(\cdot)/\partial v$.
-Moreover, the first and second fundamental forms of $f$ can be written as
-$\operatorname{I}=Edu^{2}+Gdv^{2},\quad\operatorname{II}=Ldu^{2}+Ndv^{2},$
-where
-$E=\left\langle{f_{u}},{f_{u}}\right\rangle,\quad G=\left\langle{f_{v}},{f_{v}}\right\rangle,\quad L=-\left\langle{f_{u}},{\nu_{u}}\right\rangle,\quad N=-\left\langle{f_{v}},{\nu_{v}}\right\rangle.$
-Then the Gauss and Weingarten formulas are as follows:
+Let $f\colon U\to\mathbb{R}^{3}$ be an immersion and $\nu$ its unit normal vector field defined by
+$\nu(u,v)=\frac{f_{u}\times f_{v}}{|f_{u}\times f_{v}|}(u,v),$
+where $\times$ is the cross product of $\mathbb{R}^{3}$, $(\cdot)_{u}=\partial(\cdot)/\partial u$ and $(\cdot)_{v}=\partial(\cdot)/\partial v$.
+Then we call the image of $f$ or $f$ itself a (regular) surface.
+For a surface $f\colon U\to\mathbb{R}$, we define
+(2.1) $E$ $=\left\langle{f_{u}},{f_{u}}\right\rangle,$ $F$ $=\left\langle{f_{u}},{f_{v}}\right\rangle,$ $G$ $=\left\langle{f_{v}},{f_{v}}\right\rangle,$
+$L$ $=-\left\langle{f_{u}},{\nu_{u}}\right\rangle,$ $M$ $=-\left\langle{f_{u}},{\nu_{v}}\right\rangle=-\left\langle{f_{v}},{\nu_{u}}\right\rangle,$ $N$ $=-\left\langle{f_{v}},{\nu_{v}}\right\rangle,$
+where $\left\langle{\cdot},{\cdot}\right\rangle$ is the Euclidean inner product.
+The functions $E,F,G$ are called the coefficients of the first fundamental form of $f$, and $L,M,N$ are the coefficients of the second fundamental form of $f$.
+Note that $EG-F^{2}>0$ holds on $U$.
+Let $\widehat{\operatorname{I}}$ and $\widehat{\operatorname{II}}$ be matrices given by
+$\widehat{\operatorname{I}}=\begin{pmatrix}E&F\\ F&G\end{pmatrix},\quad\widehat{\operatorname{II}}=\begin{pmatrix}L&M\\ M&N\end{pmatrix}$.
+Then $\widehat{\operatorname{I}}$ is invertible. We define the Weingarten matrix $W$ as
+$W=\widehat{\operatorname{I}}^{-1}\widehat{\operatorname{II}}=\frac{1}{EG-F^{2}}\begin{pmatrix}GL-FM&GM-FN\\ EM-FL&EN-FM\end{pmatrix}$.
+The Gaussian curvature $K$ and mean curvature $H$ of $f$ are defined by
+$K=\det{W},\quad H=\frac{1}{2}\operatorname{tr}W.$
+Moreover, the eigenvalues $\kappa_{1},\kappa_{2}$ of $W$ are called the principal curvatures of $f$. A point $p\in U$ is called an umbilic point of $f$ if $\kappa_{1}(p)=\kappa_{2}(p)$.
+Let $f\colon U\to\mathbb{R}^{3}$ be a surface and $\nu$ its unit normal vector field of $f$.
+Assume that $p\in U$ is not an umbilic point of $f$.
+Then there exist a sufficiently small neighborhood $\tilde{U}\subset U$ of $p$ and local coordinate system $(u,v)$ such that $F=M=0$ on $\tilde{U}$ (cf.
+[3, page 188]).
+We call such a coordinate system $(u,v)$ a curvature line coordinate system on $\tilde{U}$.
+In the following, we assume that a surface $f\colon U\to\mathbb{R}^{3}$ has no umbilic point and parametrized by a curvature line coordinate system $(u,v)$.
+We define $\bm{e}_{1},\bm{e}_{2}\colon U\to\mathbb{R}^{3}$ by
+(2.2) $\bm{e}_{1}=\frac{f_{u}}{|f_{u}|},\quad\bm{e}_{2}=\frac{f_{v}}{|f_{v}|}.$
+Then replacing $\nu$ with $\nu$ if necessary it holds that
+$\left\langle{\bm{e}_{i}},{\bm{e}_{j}}\right\rangle=\delta_{ij},\quad\bm{e}_{1}\times\bm{e}_{2}=\nu,$
+where $\delta_{ij}$ is the Kronecker delta. The Gauss and Weingarten formulas are as follows:
 (2.3) $f_{uu}$ $=\Gamma^{1}_{11}f_{u}+\Gamma^{2}_{11}f_{v}+L\nu=\frac{E_{u}}{2E}f_{u}-\frac{E_{v}}{2G}f_{v}+L\nu,$
@@
 $\nu_{u}$ $=-\kappa_{1}f_{u},\quad\nu_{v}=-\kappa_{2}f_{v},$
-where $\kappa_{1}$ and $\kappa_{2}$ are principal curvatures of $f$ given by
-$\kappa_{1}=\frac{L}{E},\quad\kappa_{2}=\frac{N}{G}.$
-Moreover, Codazzi equations can be written as
+where $\kappa_{1}=L/E$ and $\kappa_{2}=N/G$. Moreover, Codazzi equations can be written as
 (2.4) $L_{v}=\frac{1}{2}(\kappa_{1}+\kappa_{2})E_{v},\quad N_{u}=\frac{1}{2}(\kappa_{1}+\kappa_{2})G_{u}.$
@@
 $f_{uv}=-\frac{(\kappa_{1})_{v}}{\kappa_{1}-\kappa_{2}}f_{u}+\frac{(\kappa_{2})_{u}}{\kappa_{1}-\kappa_{2}}f_{v}.$
-Definition 2.2.
-For a surface $f\colon U\to\mathbb{R}^{3}$ parametrized by curvature line coordinate system $(u,v)$,
-  • a point $p\in U$ is said to be a $\partial_{u}$-ridge point (resp. $\partial_{v}$-ridge point) if $(\kappa_{1})_{u}(p)=0$ (resp. $(\kappa_{2})_{v}(p)=0$);
-  • a point $p\in U$ is said to be a $\partial_{u}$-sub-parabolic point (resp. $\partial_{v}$-sub-parabolic point) if $(\kappa_{2})_{u}(p)=0$ (resp. $(\kappa_{1})_{v}(p)=0$).
-Ridge points on a surface are originally introduced by Porteous to investigate singularities of distance squared functions on the surface ([20]).
-For details of ridge and sub-parabolic points on a surface, see [4, 3, 20, 21, 18].
-2.1.2. Curvature line frame
-Let $f\colon U\to\mathbb{R}^{3}$ be a regular surface without umbilic point.
-Let $\nu$ be a unit normal vector to $f$.
-We then take a curvature line coordinate system $(u,v)$ on $U$.
-We define maps $\bm{e}_{i}\colon U\to\mathbb{R}^{3}$ ($i=1,2$) by
-(2.3) $\bm{e}_{1}=\frac{f_{u}}{|f_{u}|},\quad\bm{e}_{2}=\frac{f_{v}}{|f_{v}|}.$
-Then it holds that
-$\left\langle{\bm{e}_{i}},{\bm{e}_{j}}\right\rangle=\delta_{ij},\quad\nu=\bm{e}_{1}\times\bm{e}_{2},$
-where $\delta_{ij}$ is the Kronecker delta. Thus the triplet $\{\bm{e}_{1},\bm{e}_{2},\nu\}$ forms an orthonormal frame along $f$. We call this frame the curvature line frame along $f$.
-Lemma 2.3 (cf. [5]).
-With the above condition, we have
-$\begin{pmatrix}\bm{e}_{1}\\ \bm{e}_{2}\\ \nu\end{pmatrix}_{u}=\begin{pmatrix}0&x_{1}&x_{2}\\ -x_{1}&0&0\\ -x_{2}&0&0\end{pmatrix}\begin{pmatrix}\bm{e}_{1}\\ \bm{e}_{2}\\ \nu\end{pmatrix},\quad\begin{pmatrix}\bm{e}_{1}\\ \bm{e}_{2}\\ \nu\end{pmatrix}_{v}=\begin{pmatrix}0&y_{1}&0\\ -y_{1}&0&y_{3}\\ 0&-y_{3}&0\end{pmatrix}\begin{pmatrix}\bm{e}_{1}\\ \bm{e}_{2}\\ \nu\end{pmatrix}$,
-where
-$x_{1}=\frac{(\kappa_{1})_{v}|f_{u}|}{(\kappa_{1}-\kappa_{2})|f_{v}|},\quad x_{2}=\kappa_{1}|f_{u}|,\quad y_{1}=\frac{(\kappa_{2})_{u}|f_{v}|}{(\kappa_{1}-\kappa_{2})|f_{u}|},\quad y_{3}=\kappa_{2}|f_{v}|.$
-Proof.
-For $\nu$, it holds that
-$\nu_{u}=-\kappa_{1}f_{u}=-\kappa_{1}|f_{u}|\bm{e}_{1},\quad\nu_{v}=-\kappa_{2}f_{v}=-\kappa_{2}|f_{v}|\bm{e}_{2}$
-by Weingarten formulas (2.1). By (2.3) and Lemma 2.1, $(\bm{e}_{1})_{u}$ and $(\bm{e}_{2})_{u}$ can be calculated as
-$(\bm{e}_{1})_{u}$ $=\frac{f_{uu}}{|f_{u}|}-\frac{f_{u}(|f_{u}|)_{u}}{|f_{u}|^{2}}=\frac{1}{|f_{u}|}\left(\Gamma^{1}_{11}f_{u}+\Gamma^{2}_{11}f_{v}+L\nu\right)-\Gamma^{1}_{11}\bm{e}_{1}$
-$=-\frac{E_{v}}{2G|f_{u}|}f_{v}+\frac{L}{|f_{u}|}\nu=\frac{(\kappa_{1})_{v}|f_{u}|}{(\kappa_{1}-\kappa_{2})|f_{v}|}\bm{e}_{2}+\kappa_{1}|f_{u}|\nu,$
-$(\bm{e}_{2})_{u}$ $=\frac{f_{uv}}{|f_{v}|}-\frac{f_{v}(|f_{v}|)_{u}}{|f_{v}|^{2}}=\frac{1}{|f_{v}|}\left(\Gamma^{1}_{12}f_{u}+\Gamma^{2}_{12}f_{v}\right)-\Gamma^{2}_{12}\bm{e}_{2}$
-$=\frac{E_{v}}{2E|f_{v}|}f_{u}=-\frac{(\kappa_{1})_{v}}{(\kappa_{1}-\kappa_{2})|f_{v}|}f_{u}=-\frac{(\kappa_{1})_{v}|f_{u}|}{(\kappa_{1}-\kappa_{2})|f_{v}|}\bm{e}_{1}.$
-By similar calculations, we get
-$(\bm{e}_{1})_{v}$ $=\frac{f_{uv}}{|f_{u}|}-\frac{\left\langle{f_{u}},{f_{uv}}\right\rangle}{|f_{u}|^{2}}=\frac{\Gamma^{2}_{12}}{|f_{u}|}f_{v}=\frac{G_{u}}{2G|f_{u}|}f_{v}=\frac{(\kappa_{2})_{u}|f_{v}|}{(\kappa_{1}-\kappa_{2})|f_{u}|}\bm{e}_{2},$
-$(\bm{e}_{2})_{v}$ $=\frac{f_{vv}}{|f_{v}|}-\Gamma^{2}_{22}\bm{e}_{2}=\frac{1}{|f_{v}|}\left(\Gamma^{1}_{22}f_{u}+N\nu\right)=-\frac{G_{u}}{2E|f_{v}|}f_{u}+\frac{N}{|f_{v}|}\nu$
-$=-\frac{(\kappa_{2})_{u}G}{(\kappa_{1}-\kappa_{2})E|f_{v}|}f_{u}+\kappa_{2}|f_{v}|\nu=-\frac{(\kappa_{2})_{u}|f_{v}|}{(\kappa_{1}-\kappa_{2})|f_{u}|}\bm{e}_{1}+\kappa_{2}|f_{v}|\nu.$
-Thus we have the assertions. ∎
-By this lemma, it holds that
-$(\bm{e}_{1})_{u}\times(\bm{e}_{1})_{v}=-\frac{\kappa_{1}(\kappa_{2})_{u}|f_{v}|}{\kappa_{1}-\kappa_{2}}\bm{e}_{1},\quad(\bm{e}_{2})_{u}\times(\bm{e}_{2})_{v}=-\frac{\kappa_{2}(\kappa_{1})_{v}|f_{u}|}{\kappa_{1}-\kappa_{2}}\bm{e}_{2}.$
-Therefore sets of singularities of $\bm{e}_{1}$ and $\bm{e}_{2}$ are
-(2.4) $S(\bm{e}_{1})$ $=\{q\in U\mid\kappa_{1}(q)=0\}\cup\{q\in U\mid(\kappa_{2})_{u}(q)=0\},$
-$S(\bm{e}_{2})$ $=\{q\in U\mid\kappa_{2}(q)=0\}\cup\{q\in U\mid(\kappa_{1})_{v}(q)=0\},$
-respectively.
-Since $S(\nu)=\kappa_{1}^{-1}(0)\cup\kappa_{2}^{-1}(0)$, if $p\in S(\nu)$, then $p\in S(\bm{e}_{1})$ or $p\in S(\bm{e}_{2})$.
-See [5] for singularities of frame mappings on regular surfaces.
 2.2. Fronts
-To investigate parallel surfaces below, we here recall the class of surface with certain singularities called fronts ([1, 28]).
+To investigate parallel surfaces below, we here recall the class of surfaces with certain singularities called fronts ([1, 25]).
 Let $f\colon U\to\mathbb{R}^{3}$ be a $C^{\infty}$ map, where $U\subset\mathbb{R}^{2}$ is an open set.
 Then $f$ is said to be a front if there exists a $C^{\infty}$ map $\nu\colon U\to\mathbb{S}^{2}$ such that
+  1. (1)
     $\left\langle{df_{q}(\bm{v})},{\nu(q)}\right\rangle=0$ for any $q\in U$ and $\bm{v}\in T_{q}\mathbb{R}^{2}$, and
+  2. (2)
     the pair $(f,\nu)\colon U\to\mathbb{R}^{3}\times\mathbb{S}^{2}$ gives an immersion.
@@
 Then a point $p\in U$ is said to be a singular point of $f$ if $f$ is not an immersion at $p$, that is, $\operatorname{rank}df_{p}<2$.
-We denote by $S(f)$ the set of singular points of $f$.
+We denote the set of singular points of $f$ by $S(f)$.
 We define a function $\lambda\colon U\to\mathbb{R}$ by
@@
 This curve $\gamma$ is called a singular curve for $f$ passing through $p=\gamma(0)$.
-On the other hand, for a singular point $p\in S(f)$ satisfying $\operatorname{rank}df_{p}=1$, there exists a non-zero vector field $\eta$ on $U$ such that $\eta_{p}$ generates the kernel of $df_{p}$.
+On the other hand, for a singular point $p\in S(f)$ satisfying $\operatorname{rank}df_{p}=1$, there exists a non-zero vector field $\eta$ on $V(\subset U)$ such that $\eta_{p}$ generates the kernel of $df_{p}$.
 This vector field $\eta$ is called a null vector field for $f$.
@@
 Let $f_{i}\colon(U_{i},p_{i})\to\mathbb{R}^{3}$ ($i=1,2$) be $C^{\infty}$ map germs.
-Then $f_{1}$ and $f_{2}$ are $\cal{A}$-equivalent if there exist diffeomorphism germs $\varphi\colon(U_{1},p_{1})\to(U_{2},p_{2})$ and $\Phi\colon(\mathbb{R}^{3},f_{1}(p_{1}))\to(\mathbb{R}^{3},f_{2}(p_{2}))$ such that $\Phi\circ f_{1}=f_{2}\circ\varphi$.
+Then $f_{1}$ and $f_{2}$ are $\mathcal{A}$-equivalent if there exist diffeomorphism germs $\varphi\colon(U_{1},p_{1})\to(U_{2},p_{2})$ and $\Phi\colon(\mathbb{R}^{3},f_{1}(p_{1}))\to(\mathbb{R}^{3},f_{2}(p_{2}))$ such that $\Phi\circ f_{1}=f_{2}\circ\varphi$.
 Definition 2.3.
@@
   1. (1)
-    $f$ is a cuspidal edge at $p(\in U)$ if $f$ is $\cal{A}$-equivalent to the germ $(u,v)\mapsto(u,v^{2},v^{3})$ at the origin;
+    $f$ is a cuspidal edge at $p(\in U)$ if $f$ is $\mathcal{A}$-equivalent to the germ $(u,v)\mapsto(u,v^{2},v^{3})$ at the origin;
   2. (2)
-    $f$ is a swallowtail at $p$ if $f$ is $\cal{A}$-equivalent to the germ $(u,v)\mapsto(u,4v^{3}+2uv,3v^{4}+uv^{2})$ at the origin;
+    $f$ is a swallowtail at $p$ if $f$ is $\mathcal{A}$-equivalent to the germ $(u,v)\mapsto(u,4v^{3}+2uv,3v^{4}+uv^{2})$ at the origin;
   3. (3)
-    $f$ is a cuspidal butterfly at $p$ if $f$ is $\cal{A}$-equivalent to the germ $(u,v)\mapsto(u,5v^{4}+2uv,4v^{5}+uv^{2})$ at the origin;
+    $f$ is a cuspidal butterfly at $p$ if $f$ is $\mathcal{A}$-equivalent to the germ $(u,v)\mapsto(u,5v^{4}+2uv,4v^{5}+uv^{2})$ at the origin;
   4. (4)
-    $f$ is a cuspidal lips at $p$ if $f$ is $\cal{A}$-equivalent to the germ $(u,v)\mapsto(u,2v^{3}+u^{2}v,3v^{4}+u^{2}v^{2})$ at the origin;
+    $f$ is a cuspidal lips at $p$ if $f$ is $\mathcal{A}$-equivalent to the germ $(u,v)\mapsto(u,2v^{3}+u^{2}v,3v^{4}+u^{2}v^{2})$ at the origin;
   5. (5)
-    $f$ is a cuspidal beaks at $p$ if $f$ is $\cal{A}$-equivalent to the germ $(u,v)\mapsto(u,2v^{3}-u^{2}v,3v^{4}-u^{2}v^{2})$ at the origin.
+    $f$ is a cuspidal beaks at $p$ if $f$ is $\mathcal{A}$-equivalent to the germ $(u,v)\mapsto(u,2v^{3}-u^{2}v,3v^{4}-u^{2}v^{2})$ at the origin.
 These are rank one singularities of fronts. For these singularities, the following criteria are known (see [11, 10, 9, 19, 25]).
-Fact 2.6.
+Fact 2.4.
 Let $f\colon U\to\mathbb{R}^{3}$ be a front and $p\in U$ a rank one singular point of $f$. Then
@@
 2.2.1. Limiting normal curvature and singular curvature
-Let $f\colon U\to\mathbb{R}^{3}$ be a front and $\nu$ its unit normal vector.
+Let $f\colon U\to\mathbb{R}^{3}$ be a front and $\nu$ its unit normal vector field.
 Assume that $f$ is a cuspidal edge at $p$.
@@
 Then along $\gamma$, we can define the following geometric invariants:
-(2.6) $\kappa_{\nu}(t)$ $=\frac{\left\langle{\hat{\gamma}^{\prime\prime}(t)},{\nu(\gamma(t))}\right\rangle}{|\hat{\gamma}^{\prime}(t)|^{2}}=-\frac{\left\langle{\hat{\gamma}^{\prime}(t)},{\nu^{\prime}(\gamma(t))}\right\rangle}{|\hat{\gamma}^{\prime}(t)|^{2}},$
+(2.6) $\kappa_{\nu}(t)$ $=\frac{\left\langle{\hat{\gamma}^{\prime\prime}(t)},{\nu(\gamma(t))}\right\rangle}{|\hat{\gamma}^{\prime}(t)|^{2}}=-\frac{\left\langle{\hat{\gamma}^{\prime}(t)},{\nu^{\prime}(\gamma(t))}\right\rangle}{|\hat{\gamma}^{\prime}(t)|^{2}}\quad{\left({}^{\prime}=\frac{d}{dt}\right)},$
 $\kappa_{s}(t)$ $=\operatorname{sgn}(\eta\lambda\cdot\det(\gamma^{\prime},\eta))\frac{\det(\hat{\gamma}^{\prime}(t),\hat{\gamma}^{\prime\prime}(t),\nu(\gamma(t)))}{|\hat{\gamma}^{\prime}(t)|^{3}},$
-where $\lambda$ is the signed area density function. We call $\kappa_{\nu}$ and $\kappa_{s}$ the limiting normal curvature and the singular curvature, respectively ([16, 22]).
-It is known that $\kappa_{\nu}$ relates to boundedness of the Gaussian curvature near a cuspidal edge.
-In fact, Saji, Umehara and Yamada [22] showed that the Gaussian curvature of $f$ is bounded on a sufficiently small neighborhood of $p$ if and only if $\kappa_{\nu}\equiv 0$ along a singular curvature $\gamma$ passing through $p$ (see Figure 1).
-(a) $\kappa_{\nu}=0$
-(b) $\kappa_{\nu}\neq 0$
-Figure 1. Cuspidal edges with vanishing $\kappa_{\nu}$ (left) and non-vanishing $\kappa_{\nu}$ (right).
-On the other hand, $\kappa_{s}$ corresponds to convexity or concavity of a cuspidal edge (see Figure 2).
-(a) $\kappa_{s}>0$
-(b) $\kappa_{s}<0$
-Figure 2. Cuspidal edges with positive $\kappa_{s}$ (left) and negative $\kappa_{s}$ (right).
-In particular, for the case where the Gaussian curvature $K$ is bounded near a cuspidal edge $p$, if $K$ takes non-negative values around $p$, then $\kappa_{s}$ takes non-positive values along the singular curve passing through $p$ ([22, Theorem 3.1]).
-2.3. Rational order of functions
-[…]
-Let $h_{1},h_{2}\colon(\mathbb{R}^{2},p)\to\mathbb{R}$ be function germs.
-Let $Z=h_{2}^{-1}(0)$.
-Define a functions germ $g=h_{1}/h_{2}\colon(\mathbb{R}^{2}\setminus Z,p)\to\mathbb{R}$.
-Then the rational order $\operatorname{ord}_{p}{g}$ of $g$ at $p$ is defined by
-$\operatorname{ord}_{p}{g}\coloneqq\operatorname{ord}_{p}{h_{1}}-\operatorname{ord}_{p}{h_{2}}.$
-[…]
-[…]
-Example 2.7.
-Let $h_{1},h_{2}\colon(\mathbb{R}^{2},0)\to\mathbb{R}$ be function germs given by
-$h_{1}(u,v)=uv,\quad h_{2}(u,v)=u^{2}-v^{2}.$
-[…]
-$g(u,v)=\frac{h_{1}(u,v)}{h_{2}(u,v)}=\frac{uv}{u^{2}-v^{2}}.$
-This cannot be extended as a bounded function on $(\mathbb{R}^{2},0)$. However, the rational order $\operatorname{ord}_{0}{g}$ of $g$ at 0 is
-$\operatorname{ord}_{0}{g}=2-2=0.$
-This implies that $g$ is rationally bounded at 0.
+where $\lambda$ is the signed area density function. We call $\kappa_{\nu}$ and $\kappa_{s}$ the limiting normal curvature and the singular curvature, respectively (see [13, 12, 20, 25]).
+Let $f\colon U\to\mathbb{R}^{3}$ be a front.
+Let $p\in U$ be a rank one singular point of $f$.
+Assume that $p$ is a non-degenerate singular point other than a cuspidal edge.
+Then the limiting normal curvature $\kappa_{\nu}$ of $f$ at $p$ is given by
+(2.7) $\kappa_{\nu}(p)=\lim_{t\to 0}\frac{\left\langle{\hat{\gamma}^{\prime\prime}(t)},{\nu(\gamma(t))}\right\rangle}{|\hat{\gamma}^{\prime}(t)|^{2}}=-\lim_{t\to 0}\frac{\left\langle{\hat{\gamma}^{\prime}(t)},{\hat{\nu}^{\prime}(t)}\right\rangle}{|\hat{\gamma}^{\prime}(t)|^{2}},$
+where $\hat{\gamma}(t)=f(\gamma(t))$, $\hat{\nu}(t)=\nu(\gamma(t))$ and $\gamma(t)$ is the singular curve of $f$ through $p=\gamma(0)$ ([12, Proposition 2.9]).
+On the other hand, assume that $p$ is a degenerate singular point.
+Take a local coordinate system $(u,v)$ satisfying $f_{v}(p)=0$.
+Then the limiting normal curvature $\kappa_{\nu}(p)$ of $f$ at $p$ is defined by
+(2.8) $\kappa_{\nu}(p)=\frac{\left\langle{f_{uu}(p)},{\nu(p)}\right\rangle}{|f_{u}(p)|^{2}}$
+([12, (2.2)]). In this case, $\kappa_{\nu}(p)$ does not depend on the choice of local coordinates satisfying $f_{v}(p)=0$ ([12, Proposition 2.9]).
 3. Parallel surfaces
@@
 (3.2) $f^{t}_{u}=(1-t\kappa_{1})f_{u},\quad f^{t}_{v}=(1-t\kappa_{2})f_{v}$
-by Rodrigues’ formula.
-Thus $\nu$ can be also a unit normal vector to $f^{t}$.
-Moreover, $f^{t}$ is a front for any $t\in\mathbb{R}\setminus\{0\}$ ([28]).
-By (2.1), the first and second fundamental forms of $f^{t}$ are
-(3.3) $\operatorname{I}^{t}=(1-t\kappa_{1})^{2}Edu^{2}+(1-t\kappa_{2})^{2}Gdv^{2},\quad\operatorname{II}^{t}=(1-t\kappa_{1})\kappa_{1}Edu^{2}+(1-t\kappa_{2})\kappa_{2}Gdv^{2}.$
-Therefore the Gaussian and mean curvatures for $f^{t}$ are
-(3.4) $K^{t}=\frac{\kappa_{1}\kappa_{2}}{(1-t\kappa_{1})(1-t\kappa_{2})}=\frac{K}{1-2tH+t^{2}K},\quad H^{t}=\frac{H-tK}{1-2tH+t^{2}K},$
-where $K$ and $H$ are the Gaussian and mean curvatures of $f$, respectively.
+by Weingarten formulas (2.3). Thus $\nu$ can be also a unit normal vector to $f^{t}$. Moreover, $f^{t}$ is a front for any $t\in\mathbb{R}\setminus\{0\}$ ([25]).
 For each fixed $t\in\mathbb{R}\setminus\{0\}$, the set of singular points $S(f^{t})$ of $f^{t}$ is
@@
 In addition, $\partial_{u}$ give a null vector field for $f^{t}$.
-By Fact 2.6, we have the following immediately.
-Proposition 3.1 ([8, Theorems 3.4 and 3.5]).
+By Fact 2.4, we have the following immediately.
+Proposition 3.1 ([5, Theorems 3.4 and 3.5]).
 Let $f\colon U\to\mathbb{R}^{3}$ be a regular surface without umbilic point and $f^{t}$ the parallel surface of $f$ at distance $t=1/\kappa_{1}(p)$.
@@
     $f^{t}$ is a cuspidal beaks at $p$ if and only if $(\kappa_{1})_{u}=(\kappa_{1})_{v}=0$, $(\kappa_{1})_{uu}\neq 0$ and $\det\operatorname{Hess}(\kappa_{1})<0$ at $p$.
-Corollary 3.2.
-With the same situations in Proposition 3.1, we have the following.
-  1. (1)
-    If the parallel surface $f^{t}$ is one of a cuspidal edge, a swallowtail or a cuspidal butterfly at $p$, then $\operatorname{ord}_{p}\tilde{\lambda}=1$, where $\tilde{\lambda}$ is the function as in (3.5).
-  2. (2)
-    If $f^{t}$ is either a cuspidal lips or a cuspidal beaks at $p$, then $\operatorname{ord}_{p}\tilde{\lambda}(p)=2$.
-Proof.
-We take a curvature line coordinate system $(u,v)$. We assume that $p$ is the origin 0 of $\mathbb{R}^{2}$. Then we remark that
-[…]
-hold ([10]).
-[…]
-$\tilde{\lambda}(u,v)=uh_{1}(u,v)+vh_{2}(u,v).$
-[…]
-[…]
-3.2. Behavior of the Gaussian curvature
-We consider behavior of the Gaussian curvature $K^{t}$ of the parallel surface $f^{t}$ of a regular surface $f$ at a distance $t=1/\kappa_{1}(p)$.
-When $f^{t}$ has non-degenerate singularity, the following assertion holds.
-Proposition 3.3.
-Let $f\colon U\to\mathbb{R}^{3}$ be a regular surface and $f^{t}$ the parallel surface with $t=1/\kappa_{1}(p)$.
-Suppose that $f^{t}$ is one of a cuspidal edge, a swallowtail or a cuspidal butterfly at $p$.
-Then the Gaussian curvature $K^{t}$ is rationally bounded at $p$ if and only if $\kappa_{2}$ vanishes at $p$.
-Moreover, $K^{t}$ is bounded near $p$ if and only if $\kappa_{2}$ vanishes on $S(f^{t})$.
-Proof.
-By [16, Corollary C], $K^{t}$ is rationally bounded at $p$ if and only if the Gauss map $\nu$ of $f^{t}$ is singular at $p$.
-This is equivalent to the condition that $\kappa_{2}(p)=0$ because $\kappa_{1}(p)\neq 0$.
-Thus the first assertion holds.
-Moreover, $K^{t}$ is bounded near $p$ if and only if the second fundamental form of $f^{t}$ vanishes on $S(f^{t})$ ([22, Theorem 3.1]).
-By (3.3), it holds that
-$\operatorname{II}^{t}=(1-t\kappa_{2})\kappa_{2}dv^{2}$
-on $S(f^{t})$. Thus the second assertion holds. ∎
-Proposition 3.4.
-Let $f\colon U\to\mathbb{R}^{3}$ be a regular surface whose parallel surface $f^{t}$ is a cuspidal lips or cuspidal beaks at $p$, where $t=1/\kappa_{1}(p)$.
-Then the Gaussian curvature $K^{t}$ of $f^{t}$ is rationally bounded at $p$ if and only if $\operatorname{ord}_{p}\kappa_{2}=2$.
-Moreover, $K^{t}$ is bounded near $p$ if and only if the set of parabolic points $\kappa_{2}^{-1}(0)$ contains $S(f^{t})$.
-Proof.
-By the assumption and Corollary 3.2, we see that the function $\tilde{\lambda}$ as in (3.5) satisfies $\operatorname{ord}_{p}\tilde{\lambda}=2$.
-By (3.4) and $t=1/\kappa_{1}(p)$, the Gaussian curvature $K^{t}$ of $f^{t}$ can be written as
-$K^{t}=\frac{\kappa_{1}\kappa_{2}}{t^{2}(\kappa_{1}(p)-\kappa_{1})(\kappa_{1}(p)-\kappa_{2})}=\kappa_{1}(p)^{2}\frac{\kappa_{1}\kappa_{2}}{\tilde{\lambda}(\kappa_{2}-\kappa_{1}(p))},$
-it follows that the rational order $\operatorname{ord}_{p}K^{t}$ of $K^{t}$ satisfies $\operatorname{ord}_{p}K^{t}=0$ at $p$ if and only if $\operatorname{ord}_{p}\kappa_{2}=2$.
-This implies that $K^{t}$ is rationally bounded at $p$ if and only if $\operatorname{ord}_{p}\kappa_{2}=2$.
-The boundedness is also proved.
-∎
+We remark that if $f^{t}$ ($t=1/\kappa_{1}(p)$) is a cuspidal lips, then $(\kappa_{1})_{uu}(p)\neq 0$.
+If we consider the parallel surface $f^{t}$ with $t=1/\kappa_{2}(p)$, $\tilde{\lambda}=\kappa_{2}-\kappa_{2}(p)$ is an identifier of singularities and $\partial_{v}$ is a null vector field for $f^{t}$.
+Thus one can obtain similar characterizations as Proposition 3.1.
 3.2. The limiting normal curvature
-We investigate the limiting normal curvature at a singular point of $f^{t}$.
-Proposition 3.5.
-Let $f\colon U\to\mathbb{R}^{3}$ be a regular surface without umbilic point and $\nu$ its unit normal vector.
-Suppose that the principal curvature $\kappa_{1}$ does not vanish on $U$.
-If the parallel surface $f^{t}$ with $t=1/\kappa_{1}(p)$ satisfies $\operatorname{rank}df^{t}_{p}=1$, then the limiting normal curvature $\kappa_{\nu}^{t}$ of $f^{t}$ satisfies
-$\kappa_{\nu}^{t}=\frac{\kappa_{1}\kappa_{2}}{\kappa_{1}-\kappa_{2}}$
-at $p$.
-Proof.
-We take a curvature line coordinates $(u,v)$ on $U$.
+We here consider the limiting normal curvature $\kappa_{\nu}^{t}$ of $f^{t}$ at a rank one singular point. In particular, we give the proof of Proposition 1.1.
+Proof of Proposition 1.1.
+We take a curvature line coordinates $(u,v)$.
+We show the case where $t=1/\kappa_{1}(p)$.
 We first consider the case where the parallel surface $f^{t}$ is a cuspidal edge at $p$.
@@
 $\hat{\nu}^{\prime}(v)$ $=-\kappa_{1}(c(v))f_{u}(c(v))u^{\prime}(v)-\kappa_{2}(c(v))f_{v}(c(v)),$
-where $^{\prime}=d/dv$ and $\hat{\nu}(v)=\nu(u(v),v)$. Thus we have
+where $^{\prime}=d/dv$ and $\hat{\nu}(v)=\nu(u(v),v)$. By (2.6), we have
 $\kappa_{\nu}^{t}(v)=-\frac{\left\langle{\hat{c}^{\prime}(v)},{\hat{\nu}^{\prime}(v)}\right\rangle}{|\hat{c}^{\prime}(v)|^{2}}=\frac{\kappa_{1}(c(v))\kappa_{2}(c(v))}{\kappa_{1}(c(v))-\kappa_{2}(c(v))}.$
@@
 $-\left\langle{\dot{\hat{c}}(u)},{\dot{\hat{\nu}}(u)}\right\rangle$ $=\frac{\kappa_{2}(c(u))(\kappa_{1}(c(u))-\kappa_{2}(c(u)))}{\kappa_{1}(c(u))}\dot{v}(u)^{2}G(c(u)).$
-Hence we have
+Hence by (2.7), we have
 $\kappa_{\nu}^{t}(p)=\lim_{u\to u_{0}}-\frac{\left\langle{\dot{\hat{c}}(u)},{\dot{\hat{\nu}}(u)}\right\rangle}{|\dot{\hat{c}}(u)|^{2}}=\frac{\kappa_{1}(p)\kappa_{2}(p)}{\kappa_{1}(p)-\kappa_{2}(p)}.$
@@
 $\kappa_{\nu}^{t}(p)=\frac{\left\langle{f^{t}_{vv}},{\nu(p)}\right\rangle}{|f^{t}_{v}(p)|^{2}}$
-(see [16, p.253]). By (3.2), we have
+by (2.8). By (3.2), we have
 $f^{t}_{vv}=-t(\kappa_{2})_{v}f_{v}+(1-t\kappa_{2})f_{vv}.$
@@
 $\kappa_{\nu}^{t}(p)=\frac{\kappa_{1}(p)\kappa_{2}(p)}{\kappa_{1}(p)-\kappa_{2}(p)}.$
-Hence we have the assertions. ∎
-By this proposition, we get the following.
-Corollary 3.6.
-With the same assumptions in Proposition 3.5, the limiting normal curvature $\kappa_{\nu}^{t}$ of $f^{t}$ vanishes at $p$ if and only if $\nu$ is singular at $p$.
-Moreover, $\bm{e}_{2}$ is also singular at $p$.
-Proof.
-By Proposition 3.5, $\kappa_{\nu}^{t}(p)=0$ if and only if $\kappa_{2}(p)=0$ under the assumptions.
-On the other hand, $\nu$ is singular at $p$ if and only if $\kappa_{2}(p)=0$ in this case.
-Moreover, by (2.4), $\bm{e}_{2}$ is singular at $p$ if $\kappa_{2}(p)=0$.
-Thus we get the conclusions.
-∎
+Hence we have the assertions. For the case of $t=1/\kappa_{2}(p)$, one can show in a similar way. ∎
 4. Focal surfaces
@@
 $C_{2}$).
-Thus focal surfaces are frontals in this case.
+It is known that $C_{1}$ (resp.
+$C_{2}$) is a front ([25, Theorem 7.1.1]).
 Moreover, the signed area density functions for $C_{1}$ and $C_{2}$ are
@@
 $C_{2}$).
-Moreover, by Lemma 2.3, it holds that $(\bm{e}_{1})_{u}\neq 0$ (resp.
-$(\bm{e}_{2})_{v}\neq 0$) at $p\in S(C_{1})$ (resp.
-$p\in S(C_{2})$).
-Hence $C_{1}$ (resp.
-$C_{2}$) is a front at $p\in S(C_{1})$ (resp.
-$p\in S(C_{2})$) (see [28]).
-For singularities of focal surfaces, we have the following.
+By Fact 2.4, we have the following.
+Proposition 4.1 (cf. [25, Theorem 7.1.1]).
+Let $f\colon U\to\mathbb{R}^{3}$ be a surface without umbilic point parametrized by a curvature line coordinate system $(u,v)$.
+Suppose that $\kappa_{1}$ (resp.
+$\kappa_{2}$) does not vanish on $U$.
+Then the focal surface $C_{1}$ (resp.
+$C_{2}$) is a cuspidl edge at $p\in U$ if and only if $(\kappa_{1})_{u}(p)=0$ and $(\kappa_{1})_{uu}(p)\neq 0$ (resp.
+$(\kappa_{2})_{v}(p)=0$ and $(\kappa_{2})_{vv}(p)\neq 0$).
+By Propositions 3.1 and 4.2, the following assertions hold.
 Proposition 4.2 (cf. [23]).
@@
 Let us take a curvature line coordinate system $(u,v)$.
-Assume that the parallel surface $f^{t}$ with $t=1/\kappa_{1}(p)$ of $f$ is a cuspidal edge at $p$.
-Then by Proposition 3.1, it holds that $(\kappa_{1})_{u}(p)\neq 0$.
-Thus $p\notin S(C_{1})$.
-This implies that the first assertion holds.
-Next, we assume that $f^{t}$ is one of a swallowtail, a cuspidal lips or a cuspidal beaks.
-Then by Proposition 3.1, it holds that
-$(\kappa_{1})_{u}(p)=0,\quad(\kappa_{1})_{uu}(p)\neq 0,\quad(\kappa_{1})_{v}(p)\neq 0.$
-On the other hand, since an identifier of singularities for $C_{1}$ can be taken as $\tilde{\lambda}^{C_{1}}=(\kappa_{1})_{u}$ and $\eta^{C_{1}}=\partial_{u}$ is a null vector field for $C_{1}$, we have $\eta^{C_{1}}\tilde{\lambda}^{C_{1}}=(\kappa_{1})_{uu}$, which does not vanish at $p$ by the assumption.
-Hence $C_{1}$ is a cuspidal edge at $p$ by Fact 2.6.
-Therefore we get the second assertion.
+The first assertion holds by the previous discussion.
+Thus we show the second assertion.
+We assume that $f^{t}$ is one of a swallowtail, a cuspidal lips or a cuspidal beaks.
+Then by Proposition 3.1, it holds that $(\kappa_{1})_{u}(p)=0$ and $(\kappa_{1})_{uu}(p)\neq 0$.
+Thus by Proposition 4.1, the assertion follows.
 For the case of $C_{2}$, one can show in a similar way. ∎
@@
 We investigate geometric invariants of focal surfaces at cuspidal edges.
-Let $f\colon U\to\mathbb{R}^{3}$ be a regular surface parametrized by a curvature line coordinate system $(u,v)$.
-We then assume that $\kappa_{2}(p)=0$ and the focal surface $C_{1}$ is a cuspidal edge at $p\in U$.
-In this case, we shall consider geometric invariants called the limiting normal curvature $\kappa_{\nu}^{C_{1}}$ and the singular curvature $\kappa_{s}^{C_{1}}$ of $C_{1}$ at $p$.
-By the assumption, $S(C_{1})=((\kappa_{1})_{u})^{-1}(0)$ and $(\kappa_{1})_{uu}(p)\neq 0$ hold.
+Let $f\colon U\to\mathbb{R}^{3}$ be a regular surface without umbilic points.
+Assume that $C_{i}$ ($i=1$ or $2$) is a cuspidal edge at $p\in U$.
+Then we consider the limiting normal curvature $\kappa_{\nu}^{C_{i}}$ and the singular curvature $\kappa_{s}^{C_{i}}$ of $C_{i}$ at $p$.
+In particular, we shall give the proof of Theorem 1.2.
+Using a curvature line coordinate system $(u,v)$ for $f$, the limiting normal curvature and the singular curvature for $C_{i}$ obtained as follows.
+Proposition 4.4.
+Let $f\colon U\to\mathbb{R}^{3}$ be a regular surface without umbilic points. Assume that $f$ is parametrized by a curvature line coordinate system $(u,v)$.
+  1. (1)
+    If $C_{1}$ is a cuspidal edge at $p\in U$, then the limiting normal curvature $\kappa_{\nu}^{C_{1}}$ and the singular curvature $\kappa_{s}^{C_{1}}$ can be written as
+    (4.4) $\kappa_{\nu}^{C_{1}}$ $=-\frac{\kappa_{1}^{3}(\kappa_{2})_{u}G}{\sqrt{E}((\kappa_{1}-\kappa_{2})^{2}\kappa_{1}^{2}G+(\kappa_{1})_{v}^{2})},$
+    $\kappa_{s}^{C_{1}}$ $=-\operatorname{sgn}((\kappa_{1})_{uu}\kappa_{1}(\kappa_{1}-\kappa_{2}))\frac{\kappa_{1}^{3}\sqrt{G}X_{1}}{((\kappa_{1}-\kappa_{2})^{2}\kappa_{1}^{2}G+(\kappa_{1})_{v}^{2})^{3/2}}$
+    at $p$, where $X_{1}$ is
+    $X_{1}=(\kappa_{1}-\kappa_{2})\left((\kappa_{1}-\kappa_{2})\kappa_{1}\kappa_{2}G-\frac{\det\operatorname{Hess}(\kappa_{1})}{(\kappa_{1})_{uu}}\right)+(\kappa_{1})_{v}(2(\kappa_{1})_{v}+\Gamma^{2}_{22}(\kappa_{1}-\kappa_{2})-(\kappa_{2})_{v}).$
+  2. (2)
+    If $C_{2}$ is a cuspidal edge at $p\in U$, then the limiting normal curvature $\kappa_{\nu}^{C_{2}}$ and the singular curvature $\kappa_{s}^{C_{2}}$ can be written as
+    (4.5) $\kappa_{\nu}^{C_{2}}$ $=-\frac{\kappa_{2}^{3}(\kappa_{1})_{v}E}{\sqrt{G}((\kappa_{1}-\kappa_{2})^{2}\kappa_{2}^{2}E+(\kappa_{2})_{u}^{2})},$
+    $\kappa_{s}^{C_{2}}$ $=-\operatorname{sgn}((\kappa_{2})_{vv}\kappa_{2}(\kappa_{2}-\kappa_{1}))\frac{\kappa_{2}^{3}\sqrt{E}X_{2}}{((\kappa_{1}-\kappa_{2})^{2}\kappa_{2}^{2}E+(\kappa_{2})_{u}^{2})^{3/2}}$
+    at $p$, where $X_{2}$ is
+    $X_{2}=(\kappa_{2}-\kappa_{1})\left((\kappa_{2}-\kappa_{1})\kappa_{1}\kappa_{2}E-\frac{\det\operatorname{Hess}(\kappa_{2})}{(\kappa_{2})_{vv}}\right)+(\kappa_{2})_{u}(2(\kappa_{2})_{u}+\Gamma^{1}_{11}(\kappa_{2}-\kappa_{1})-(\kappa_{1})_{u}).$
+Proof.
+We show the case of $C_{1}$.
+By (4.3) and Proposition 4.1, $S(C_{1})=((\kappa_{1})_{u})^{-1}(0)$ and $(\kappa_{1})_{uu}(p)\neq 0$ hold.
 Thus there exists a regular curve $\gamma(v)=(u(v),v)$ ($|v-v_{0}|<\varepsilon$) with $\gamma(v_{0})=p$ such that $(\kappa_{1})_{u}(u(v),v)=0$ by the implicit function theorem.
-The curve $\gamma(v)$ is a singular curve for $C_{1}$.
-Thus the curve
+The curve $\gamma(v)$ is a singular curve for $C_{1}$, and the curve
 (4.6) $\hat{\gamma}(v)=C_{1}(\gamma(v))=C_{1}(u(v),v)$
-gives the singular locus of $C_{1}$ near $p$. We investigate the first and second differentials of $\hat{\gamma}$. By (4.2) and $(\kappa_{1})_{u}(u(v),v)=0$, it holds that
+gives the singular locus of $C_{1}$ near $p$. By (4.2) and $(\kappa_{1})_{u}(u(v),v)=0$, it holds that
 (4.7) $\hat{\gamma}^{\prime}(v)$ $=(C_{1})_{v}(\gamma(v))=\left(1-\frac{\kappa_{2}(\gamma(v))}{\kappa_{1}(\gamma(v))}\right)f_{v}(\gamma(v))-\frac{(\kappa_{1})_{v}(\gamma(v))}{\kappa_{1}(\gamma(v))^{2}}\nu(\gamma(v)),$
@@
 $(C_{1})_{vv}$ $=\left(1-\frac{\kappa_{2}}{\kappa_{1}}\right)f_{vv}+\left(\frac{2\kappa_{2}(\kappa_{1})_{v}}{\kappa_{1}^{2}}-\frac{(\kappa_{2})_{v}}{\kappa_{1}}\right)f_{v}+\left(\frac{2(\kappa_{1})_{v}^{2}}{\kappa_{1}^{3}}-\frac{(\kappa_{1})_{vv}}{\kappa_{1}^{2}}\right)\nu$
-$=\left(1-\frac{\kappa_{2}}{\kappa_{1}}\right)\Gamma^{1}_{22}f_{u}+\left(\frac{2\kappa_{2}(\kappa_{1})_{v}}{\kappa_{1}^{2}}-\frac{(\kappa_{2})_{v}}{\kappa_{1}}+\left(1-\frac{\kappa_{2}}{\kappa_{1}}\right)\Gamma^{2}_{22}\right)f_{v}$
-$\hskip 142.26378pt+\left(\frac{2(\kappa_{1})_{v}^{2}}{\kappa_{1}^{3}}-\frac{(\kappa_{1})_{vv}}{\kappa_{1}^{2}}+\left(1-\frac{\kappa_{2}}{\kappa_{1}}\right)N\right)\nu$
+$=\left(1-\frac{\kappa_{2}}{\kappa_{1}}\right)\Gamma^{1}_{22}f_{u}+\left(\frac{2\kappa_{2}(\kappa_{1})_{v}}{\kappa_{1}^{2}}-\frac{(\kappa_{2})_{v}}{\kappa_{1}}+\left(1-\frac{\kappa_{2}}{\kappa_{1}}\right)\Gamma^{2}_{22}\right)f_{v}+\left(\frac{2(\kappa_{1})_{v}^{2}}{\kappa_{1}^{3}}-\frac{(\kappa_{1})_{vv}}{\kappa_{1}^{2}}+\left(1-\frac{\kappa_{2}}{\kappa_{1}}\right)N\right)\nu$
 along $\gamma$. On the other hand, differentiating $(\kappa_{1})_{u}(u(v),v)=0$ by $v$, we get
@@
 near $p$. By (4.7), (4.8) and (4.9), the second derivative of $\hat{\gamma}$ is
-(4.8) $\hat{\gamma}^{\prime\prime}(v)=\left(1-\frac{\kappa_{2}(\gamma(v))}{\kappa_{1}(\gamma(v))}\right)\Gamma^{1}_{22}(\gamma(v))f_{u}(\gamma(v))+\left(\frac{2\kappa_{2}(\gamma(v))(\kappa_{1})_{v}(\gamma(v))}{\kappa_{1}(\gamma(v))^{2}}-\frac{(\kappa_{2})_{v}(\gamma(v))}{\kappa_{1}(\gamma(v))}+\left(1-\frac{\kappa_{2}(\gamma(v))}{\kappa_{1}(\gamma(v))}\right)\Gamma^{2}_{22}(\gamma(v))\right)f_{v}(\gamma(v))+\left(-\frac{\det\operatorname{Hess}(\kappa_{1})(\gamma(v))}{\kappa_{1}(\gamma(v))^{2}(\kappa_{1})_{uu}(\gamma(v))}+\frac{2(\kappa_{1})_{v}(\gamma(v))^{2}}{\kappa_{1}(\gamma(v))^{3}}+\left(1-\frac{\kappa_{2}(\gamma(v))}{\kappa_{1}(\gamma(v))}\right)N(\gamma(v))\right)\nu(\gamma(v)),$
-where $\det\operatorname{Hess}(\kappa_{1})(\gamma(v))=(\kappa_{1})_{uu}(\gamma(v))(\kappa_{1})_{vv}(\gamma(v))-(\kappa_{1})_{uv}(\gamma(v))^{2}$.
-For the limiting normal curvature, we get the following assertion.
-Proposition 4.3.
-Let $f\colon U\to\mathbb{R}^{3}$ be a regular surface parametrized by a curvature line coordinates $(u,v)$.
-Suppose that the focal surface $C_{1}$ of $f$ is a cuspidal edge at $p$ and $\kappa_{2}(p)=0$.
-Then the limiting normal curvature $\kappa_{\nu}^{C_{1}}$ at $p$ is calculated as
-$\kappa_{\nu}^{C_{1}}(p)=-\frac{\kappa_{1}(p)^{3}(\kappa_{2})_{u}(p)G(p)}{\sqrt{E(p)}\left(\kappa_{1}(p)^{4}G(p)+(\kappa_{1})_{v}(p)^{2}\right)}$
-Proof.
+(4.10) $\hat{\gamma}^{\prime\prime}(v)$ $=\left(1-\frac{\kappa_{2}(\gamma(v))}{\kappa_{1}(\gamma(v))}\right)\Gamma^{1}_{22}(\gamma(v))f_{u}(\gamma(v))$
+$\quad+\left(\frac{2\kappa_{2}(\gamma(v))(\kappa_{1})_{v}(\gamma(v))}{\kappa_{1}(\gamma(v))^{2}}-\frac{(\kappa_{2})_{v}(\gamma(v))}{\kappa_{1}(\gamma(v))}+\left(1-\frac{\kappa_{2}(\gamma(v))}{\kappa_{1}(\gamma(v))}\right)\Gamma^{2}_{22}(\gamma(v))\right)f_{v}(\gamma(v))$
+$\quad\quad+\left(-\frac{\det\operatorname{Hess}(\kappa_{1})(\gamma(v))}{\kappa_{1}(\gamma(v))^{2}(\kappa_{1})_{uu}(\gamma(v))}+\frac{2(\kappa_{1})_{v}(\gamma(v))^{2}}{\kappa_{1}(\gamma(v))^{3}}+\left(1-\frac{\kappa_{2}(\gamma(v))}{\kappa_{1}(\gamma(v))}\right)N(\gamma(v))\right)\nu(\gamma(v)).$
 Since $\bm{e}_{1}$ is a unit normal vector to $C_{1}$, the limiting normal curvature of $C_{1}$ at $p$ is
 $\kappa_{\nu}^{C_{1}}(p)=\frac{\left\langle{\hat{\gamma}^{\prime\prime}(v_{0})},{\bm{e}_{1}(p)}\right\rangle}{|\hat{\gamma}^{\prime}(v_{0})|^{2}}$
-by (2.6), where $\gamma(v)$ is a singular curve for $C_{1}$ with $\gamma(v_{0})=p$ and $\hat{\gamma}=C_{1}\circ\gamma$. By (2.1), (4.5), (4.8) and Lemma 2.1, we have
-$|\hat{\gamma}^{\prime}(v_{0})|^{2}$ $=\frac{1}{\kappa_{1}(p)^{4}}\left(\kappa_{1}(p)^{4}G(p)+(\kappa_{1})_{v}(p)^{2}\right),$
-$\left\langle{\hat{\gamma}^{\prime\prime}(v_{0})},{\bm{e}_{1}(p)}\right\rangle$ $=\Gamma^{1}_{22}(p)\sqrt{E(p)}=-\frac{G_{u}}{2\sqrt{E(p)}}=-\frac{(\kappa_{2})_{u}(p)G(p)}{\kappa_{1}(p)\sqrt{E(p)}}.$
-Thus we obtain the assertion. ∎
-Corollary 4.4.
-With the same situations in Proposition 4.3, the point $p$ is a $\partial_{u}$-subparabolic point for $f$ if and only if $\kappa_{\nu}^{C_{1}}$ vanishes at $p$.
-Moreover, $\bm{e}_{1}$ is singular at $p$ when $\kappa_{\nu}^{C_{1}}(p)=0$.
-Furthermore, the Gaussian curvature $K^{C_{1}}$ of $C_{1}$ is rationally bounded at $p$ in that case.
-Proof.
-The first assertion follows from Proposition 4.3. We get the second statement by Lemma 2.3. The last assertion holds by (A.6) and [16, Corollary C]. ∎
-For the singular curvature $\kappa_{s}^{C_{1}}$ of $C_{1}$, we have the following.
-Proposition 4.5.
-Let $f\colon U\to\mathbb{R}^{3}$ be a regular surface parametrized by curvature line coordinates $(u,v)$.
-Suppose that $\kappa_{2}(p)=0$ and the focal surface $C_{1}$ is a cuspidal edge at $p$.
-Then the singular curvature $\kappa_{s}^{C_{1}}$ of $C_{1}$ is
-$\kappa_{s}^{C_{1}}=\operatorname{sgn}((\kappa_{1})_{uu})\frac{\kappa_{1}^{3}\sqrt{G}}{(\kappa_{1})_{uu}\delta^{3}}\left(\kappa_{1}\det\operatorname{Hess}(\kappa_{1})-(\kappa_{1})_{v}(\kappa_{1})_{uu}\left(\Gamma^{2}_{22}\kappa_{1}+2(\kappa_{1})_{v}-(\kappa_{2})_{v}\right)\right)$
-at $p$, where we set
-$\delta=\sqrt{\kappa_{1}(p)^{4}G(p)+(\kappa_{1})_{v}(p)^{2}}.$
-Proof.
-Since $\bm{e}_{1}$ is a unit normal vector to $C_{1}$, the singular curvature $\kappa_{s}^{C_{1}}$ at $p$ is
-$\kappa_{s}^{C_{1}}(p)=\operatorname{sgn}(\det(\gamma^{\prime},\eta)(v_{0})\cdot\eta^{C_{1}}\lambda^{C_{1}}(p))\frac{\det(\hat{\gamma}^{\prime}(v_{0}),\hat{\gamma}^{\prime\prime}(v_{0}),\bm{e}_{1}(p))}{|\hat{\gamma}^{\prime}(v_{0})|^{3}}.$
+by (2.6). By (2.3), (4.7), (4.10) and Lemma 2.1, we have
+$|\hat{\gamma}^{\prime}(v_{0})|^{2}$ $=\frac{1}{\kappa_{1}(p)^{4}}\left((\kappa_{1}(p)-\kappa_{2}(p))^{2}\kappa_{1}(p)^{2}G(p)+(\kappa_{1})_{v}(p)^{2}\right),$
+$\left\langle{\hat{\gamma}^{\prime\prime}(v_{0})},{\bm{e}_{1}(p)}\right\rangle$ $=\frac{\kappa_{1}(p)-\kappa_{2}(p)}{\kappa_{1}(p)}\Gamma^{1}_{22}(p)\sqrt{E(p)}=-\frac{\kappa_{1}(p)-\kappa_{2}(p)}{\kappa_{1}(p)}\frac{G_{u}}{2\sqrt{E(p)}}=-\frac{(\kappa_{2})_{u}(p)G(p)}{\kappa_{1}(p)\sqrt{E(p)}}.$
+Thus we obtain the assertion for $\kappa_{\nu}^{C_{1}}$.
+We next consider $\kappa_{s}^{C_{1}}$. The singular curvature $\kappa_{s}^{C_{1}}$ at $p$ is
+$\kappa_{s}^{C_{1}}(p)=\operatorname{sgn}(\det(\gamma^{\prime},\eta^{C_{1}})(v_{0})\cdot\eta^{C_{1}}\lambda^{C_{1}}(p))\frac{\det(\hat{\gamma}^{\prime}(v_{0}),\hat{\gamma}^{\prime\prime}(v_{0}),\bm{e}_{1}(p))}{|\hat{\gamma}^{\prime}(v_{0})|^{3}}.$
 By the scalar triple product, it holds that
@@
 By (4.7), we see that
-$\bm{e}_{1}(p)\times\hat{\gamma}^{\prime}(v_{0})=\frac{(\kappa_{1})_{v}(p)}{\kappa_{1}(p)^{2}}\bm{e}_{2}(p)+|f_{v}(p)|\nu(p).$
+$\bm{e}_{1}(p)\times\hat{\gamma}^{\prime}(v_{0})=\frac{(\kappa_{1})_{v}(p)}{\kappa_{1}(p)^{2}}\bm{e}_{2}(p)+\frac{\kappa_{1}(p)-\kappa_{2}(p)}{\kappa_{1}(p)}\sqrt{G(p)}\nu(p).$
 Thus by (4.10), we have
-$\det(\hat{\gamma}^{\prime}(v_{0}),\hat{\gamma}^{\prime\prime}(v_{0}),\bm{e}_{1}(p))$
-$=\frac{-|f_{v}(p)|}{\kappa_{1}(p)^{3}(\kappa_{1})_{uu}(p)}\Big{(}\kappa_{1}(p)\det\operatorname{Hess}(\kappa_{1})(p)$
-$\qquad\qquad-(\kappa_{1})_{v}(p)(\kappa_{1})_{uu}(p)\left(\kappa_{1}(p)\Gamma^{2}_{22}(p)+2(\kappa_{1})_{v}(p)-(\kappa_{2})_{v}(p)\right)\Big{)}.$
+$\det(\hat{\gamma}^{\prime}(v_{0}),\hat{\gamma}^{\prime\prime}(v_{0}),\bm{e}_{1}(p))=\left\langle{\bm{e}_{1}(p)\times\hat{\gamma}^{\prime}(v_{0})},{\hat{\gamma}^{\prime\prime}(v_{0})}\right\rangle=\frac{\sqrt{G}(p)}{\kappa_{1}(p)^{3}}X_{1}(p).$
 On the other hand, since $\eta^{C_{1}}=\partial_{u}$, we get
-$\eta^{C_{1}}\lambda^{C_{1}}(p)=\frac{(\kappa_{1})_{uu}(p)}{\kappa_{1}(p)^{2}}|f_{v}(p)|,\quad\det(\gamma^{\prime},\eta)(v_{0})=-1$
+$\eta^{C_{1}}\lambda^{C_{1}}(p)=\frac{(\kappa_{1})_{uu}(p)(\kappa_{1}(p)-\kappa_{2}(p))}{\kappa_{1}(p)^{3}}|f_{v}(p)|,\quad\det(\gamma^{\prime},\eta)(v_{0})=-1$
 by (4.3). Hence it follows that
-$\operatorname{sgn}(\det(\gamma^{\prime},\eta)(v_{0})\cdot\eta^{C_{1}}\lambda^{C_{1}}(p))=-\operatorname{sgn}((\kappa_{1})_{uu}(p)).$
-Since
-$|\hat{\gamma}^{\prime}(v_{0})|=\frac{1}{\kappa_{1}(p)^{2}}\sqrt{(\kappa_{1})(p)^{4}G(p)+(\kappa_{1})_{v}(p)^{2}}=\frac{\delta}{\kappa_{1}(p)^{2}},$
-we have the conclusion. ∎
-By the expression as in Proposition 4.5, when the parallel surface $f^{t}$ with $t=1/\kappa_{1}(p)$ is a swallowtail at $p$, $\kappa_{s}^{C_{1}}$ takes either a positive or negative value at $p$.
-However, when $f^{t}$ is either a cuspidal lips or a cuspidal beaks at $p$, one can notice the following.
-Theorem 4.6.
-Let $f\colon U\to\mathbb{R}^{3}$ be a regular surface whose parallel surface $f^{t}$ is a cuspidal lips (resp.
-a cuspidal beaks) at $p\in U$, where $t=1/\kappa_{1}(p)$.
-Suppose that $\kappa_{2}(p)=0$.
-Then the singular curvature $\kappa_{s}^{C_{1}}$ at a cuspidal edge $p$ of $C_{1}$ is positive (resp.
-negative).
-Proof.
+$\operatorname{sgn}(\det(\gamma^{\prime},\eta)(v_{0})\cdot\eta^{C_{1}}\lambda^{C_{1}}(p))=-\operatorname{sgn}((\kappa_{1})_{uu}(p)\kappa_{1}(p)(\kappa_{1}(p)-\kappa_{2}(p))).$
+Thus we have the conclusion for $\kappa_{s}^{C_{1}}$.
+For $C_{2}$, one can show in a similar manner. ∎
+By using Proposition 4.4, we show Theorem 1.2.
+Proof of Theorem 1.2.
+We show the case of $C_{1}$.
+By Proposition 1.1, $\kappa_{2}(p)=0$.
 Let us take a curvature line coordinate system $(u,v)$.
@@
 Therefore $\operatorname{sgn}(\kappa_{s}^{C_{1}}(p))=\operatorname{sgn}(\det\operatorname{Hess}(\kappa_{1})(p))$ holds. Hence we get the assertion. ∎
-If the Gaussian curvature $K^{t}$ of $f^{t}$ with $t=1/\kappa_{1}(p)$ is at least ratinally bounded at $p$, then $\kappa_{2}(p)=0$ by Proposition 3.4. Thus we get the following.
-Corollary 4.7.
-Let $f\colon U\to\mathbb{R}^{3}$ be a regular surface whose parallel surface $f^{t}$ is a cuspidal lips (resp. cuspidal beaks) at $p\in U$, where $t=1/\kappa_{1}(p)$.
-  1. (1)
-    If the Gaussian curvature $K^{t}$ of $f^{t}$ is rationally bounded at $p$, then the singular curvature $\kappa_{s}^{C_{1}}$ for the focal surface $C_{1}$ of $f$ is positive (resp.
-negative) at $p$.
-  2. (2)
-    If the Gaussian curvature $K^{t}$ of $f^{t}$ is bounded near $p$, then $\kappa_{s}^{C_{1}}$ for $C_{1}$ is positive (resp. negative) at $p$.
 Example 4.5.
@@
 Moreover, we see that the parallel surface $f^{1}=f+\nu$ is a cuspidal beaks at $(0,0)$ (see Figure 3).
-We denote the coefficients of the first and second fundamental forms by $E,F,G,L,M$ and $N$.
-Note that $N=0$.
-Then one can take a principal vector $\bm{v}_{1}$ with respect to $\kappa_{1}$ is
-$\bm{v}_{1}=(\kappa_{1}G,M-\kappa_{1}F).$
-This is a null vector of $C_{1}$.
-It is known that $(\bm{v}_{1}\kappa_{1})^{-1}(0)$ corresponds to the set of singular points of $C_{1}$ ([20]).
-We remark that $((\bm{v}_{1}\kappa_{1})_{u},(\bm{v}_{1}\kappa_{1})_{v})=(21,6)\neq(0,0)$ at $(0,0)$.
-Thus a vector field
-$\xi=-(\bm{v}_{1}\kappa_{1})_{v}\partial_{u}+(\bm{v}_{1}\kappa_{1})_{u}\partial_{v}$
-is tangent to $(\bm{v}_{1}\kappa_{1})^{-1}(0)$.
-Then it follows that $\det(\xi{C_{1}},\xi\xi{C_{1}},\bm{e}_{1})(0,0)=15876>0$, where $\bm{e}_{1}=\bm{v}_{1}f/|\bm{v}_{1}f|$ is a unit normal vector to $C_{1}$.
-On the other hand, it holds that
-$\bm{v}_{1}\lambda^{C_{1}}=21\quad\det(\xi,\bm{v}_{1})=-21$
-at $(0,0)$. Thus we have
-$\kappa_{s}^{C_{1}}(0)=\operatorname{sgn}(\bm{v}_{1}\lambda^{C_{1}}(0,0)\cdot\det(\xi,\bm{v}_{1})(0,0))\frac{\det(\xi{C_{1}},\xi\xi{C_{1}},\bm{e}_{1})}{|\xi{C_{1}}|^{3}}(0,0)=-\frac{12}{7}<0.$
+We find that the limiting normal curvature of $f^{1}$ vanishes at $(0,0)$.
+In addition, the singular curvature $\kappa_{s}^{C_{1}}$ of $C_{1}$ is $\kappa_{s}^{C_{1}}(0)=-12/7<0$.
 Figure 3. The images of $f$ (blue), $f^{1}$ (purple) and $C_{1}$ (gray).
@@
 Let $f\colon\mathbb{R}^{2}\to\mathbb{R}^{3}$ be a regular surface given by
-$f(u,v)=\left(u,v,\frac{1}{2}u+uv^{2}+u^{4}\right).$
+$f(u,v)=\left(u,v,\frac{1}{2}u^{2}+uv^{2}+u^{4}\right).$
 Then
@@
 By direct calculation, we see that $f^{1}=f+\nu$ is a cuspidal lips at $(0,0)$.
-Moreover, following similar calculations as in the previous example, it can be observed that the singular curvature $\kappa_{s}^{C_{1}}$ of the focal surface $C_{1}$ is positive at $(0,0)$ (see Figure 4).
-Indeed, we have $\kappa_{s}^{C_{1}}(0)=8>0$.
+Moreover, it can be observed that the singular curvature $\kappa_{s}^{C_{1}}$ of the focal surface $C_{1}$ is $\kappa_{s}^{C_{1}}(0)=8>0$ at $(0,0)$ (see Figure 4).
 Figure 4. The images of $f$ (red), $f^{1}$ (green) and $C_{1}$ (purple).
-Appendix A Gaussian curvature of the focal surface
-For a surface $f\colon U\to\mathbb{R}^{3}$ parametrized by curvature line coordinates $(u,v)$, we recall the formula for the Gaussian curvature of the focal surface $C_{1}$.
-By (4.2), the first fundamental form of $C_{1}$ is
-(A.1) $\operatorname{I}^{C_{1}}=\frac{(\kappa_{1})_{u}^{2}}{\kappa_{1}^{4}}du^{2}+\frac{2(\kappa_{1})_{u}(\kappa_{1})_{v}}{\kappa_{1}^{4}}dudv+\left(\left(1-\frac{\kappa_{2}}{\kappa_{1}}\right)^{2}G+\frac{(\kappa_{1})_{v}^{2}}{\kappa_{1}^{4}}\right)dv^{2}.$
-Thus we get
-(A.2) $\det\operatorname{I}^{C_{1}}=\frac{(\kappa_{1})_{u}^{2}}{\kappa_{1}^{4}}\left(1-\frac{\kappa_{2}}{\kappa_{1}}\right)^{2}G=\frac{(\kappa_{1})_{u}^{2}(\kappa_{1}-\kappa_{2})^{2}G}{\kappa_{1}^{6}}.$
-On the other hand, the second fundamental form $\operatorname{II}^{C_{1}}$ of $C_{1}$ with respect to $\bm{n}$ is given by
-(A.3) $\operatorname{II}^{C_{1}}=-\frac{(\kappa_{1})_{u}L}{\kappa_{1}^{2}|f_{u}|}du^{2}+\frac{G_{u}}{2|f_{u}|}\left(1-\frac{\kappa_{2}}{\kappa_{1}}\right)dv^{2}.$
-By Lemma 2.1, we see that
-$\frac{G_{u}}{2}=\frac{(\kappa_{2})_{u}G}{\kappa_{1}-\kappa_{2}}.$
-Substituting this for (A.3), we obtain
-(A.4) $\operatorname{II}^{C_{1}}=-\frac{(\kappa_{1})_{u}L}{\kappa_{1}^{2}|f_{u}|}du^{2}+\frac{(\kappa_{2})_{u}G}{\kappa_{1}|f_{u}|}dv^{2}.$
-Thus we have
-(A.5) $\det\operatorname{II}^{C_{1}}=-\frac{(\kappa_{1})_{u}(\kappa_{2})_{u}GL}{E\kappa_{1}^{3}}=-\frac{(\kappa_{1})_{u}(\kappa_{2})_{u}G}{\kappa_{1}^{2}}.$
-By (A.2) and (A.5), the Gaussian curvature $K^{C_{1}}$ of $C_{1}$ can be calculated as
-(A.6) $K^{C_{1}}=\frac{\det\operatorname{II}^{C_{1}}}{\det\operatorname{I}^{C_{1}}}=-\frac{\kappa_{1}^{4}(\kappa_{2})_{u}}{(\kappa_{1})_{u}(\kappa_{1}-\kappa_{2})^{2}}.$
+Example 4.7.
+Let $f\colon\mathbb{R}^{2}\to\mathbb{R}^{3}$ be a regular surface given by
+$f(u,v)=\left(u,v,\frac{1}{2}u^{2}-v^{2}+u^{4}+u^{3}v\right).$
+Then the unit normal vector field $\nu$ is
+$\nu(u,v)=\frac{(-u(1+4u^{2}+3uv),-u^{3}+2v,1)}{\sqrt{1+(u^{3}+2v)^{2}+u^{2}(1+4u^{2}+3uv)^{2}}}.$
+Principal curvatures $\kappa_{1},\kappa_{2}$ of $f$ can be chosen as
+$\kappa_{1}(0,0)=-2,\quad\kappa_{2}(0,0)=1\neq 0.$
+Then a parallel surface $f^{-1/2}=f-\nu/2$ has a cuspidal lips at $(0,0)$.
+Moreover, the singular curvature of $C_{1}$ at the origin is calculated as $\kappa_{s}^{C_{1}}(0)=-9/4<0$ (see Figure 5) This gives a counterexample of Theorem 1.2.
+Figure 5. The image of $f^{-1/2}$ (left), $C_{1}$ (middle) and both of them (right).
+If the Gaussian curvature of $f^{t}$ is bounded near its singular point $p$, then $\kappa_{\nu}^{t}$ vanishes at $p$ (see Appendix A). Thus we get the following by Theorem 1.2 immediately.
+Corollary 4.8.
+Let $f\colon U\to\mathbb{R}^{3}$ be a regular surface whose parallel surface $f^{t}$ is a cuspidal lips (resp.
+cuspidal beaks) at $p\in U$, where $t=1/\kappa_{i}(p)$ ($i=1$ or $2$).
+If the Gaussian curvature $K^{t}$ of $f^{t}$ is bounded near $p$, then $\kappa_{s}^{C_{i}}$ for $C_{i}$ is positive (resp.
+negative) at $p$.
+Appendix A Gaussian curvature of parallels and focal surfaces
+Let $f\colon U\to\mathbb{R}^{3}$ be a regular surface and $\nu$ its unit normal vector field of $f$. Then the Gaussian curvature $K^{t}$ of the parallel surface $f^{t}=f+t\nu$ is
+$K^{t}=\frac{\kappa_{1}\kappa_{2}}{(1-t\kappa_{1})(1-t\kappa_{2})}$
+(cf.
+[4]).
+Thus when $t=1/\kappa_{1}(p)$ (resp.
+$t=1/\kappa_{2}(p)$), $K^{t}$ is bounded near $p$ if and only if $\kappa_{1}^{-1}(\kappa_{1}(p))\subset\kappa_{2}^{-1}(0)$ (resp.
+$\kappa_{2}^{-1}(\kappa_{2}(p))\subset\kappa_{1}^{-1}(0)$) (cf.
+[20, 12]).
+Moreover, the Gaussian curvature $K^{C_{i}}$ ($i=1,2$) of $C_{i}$ can be calculated as
+(A.1) $K^{C_{i}}=\begin{dcases}-\frac{\kappa_{1}^{4}(\kappa_{2})_{u}}{(\kappa_{1})_{u}(\kappa_{1}-\kappa_{2})^{2}}&(i=1),\\ -\frac{\kappa_{2}^{4}(\kappa_{1})_{v}}{(\kappa_{2})_{v}(\kappa_{1}-\kappa_{2})^{2}}&(i=2)\end{dcases}$
+([15, 4]).
 Thus the Gaussian curvature $K^{C_{1}}$ of $C_{1}$ is bounded near a singular point $p$ if and only if $((\kappa_{1})_{u})^{-1}(0)\subset((\kappa_{2})_{u})^{-1}(0)$.